### v2

# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

Abstract
[…]
1 Introduction
Diffusion models have become the de-facto standard for high-quality image generation, owing to their strong scalability with data, model size, and compute. This scalability has driven remarkable advances in text-to-image synthesis [ho2020denoising, saharia2022photorealistic, rombach2022high, ruiz2023dreambooth, sun2024coma, sun2025ouroboros], establishing scaling laws as a guiding principle for building more capable models. While most prior work has focused on scaling at training time, recent studies show that test-time scaling (TTS), allocating additional compute during inference, can also significantly improve sample quality and overall performance [nichol2021glide, esser2024scaling, peebles2023scalable, ma2025scaling, liu2024correcting]. Despite its promise, existing TTS research remains limited in scope, leaving open questions about how to use inference-time compute more effectively.
Figure 1: Global vs. Local Test-Time Scaling. Conventional TTS methods perform global search, sampling or perturbing the entire image, which ignores the inherent spatial heterogeneity of image quality and wastes computation on regions that are already good. LoTTS instead performs localized refinement: it identifies defective regions using quality-aware masks and selectively resamples only where needed, improving low-quality areas while preserving high-quality content.
Existing TTS methods can be broadly grouped into three categories. The first is denoising step scaling, which improves quality by increasing the number of sampling steps [song2020score, lu2022dpm]. However, these gains saturate quickly and plateau around 50 steps, with further increases offering negligible benefit. The second is Best-of-$N$ search, where generates $N$ samples and selects the best one by verifier [liu2024correcting, wang2023exploring]. While simple, this brute-force approach treats each candidate as an independent sample from scratch, overlooking the fact that even imperfect images may be substantially improved through local corrections. As a result, potentially promising samples are discarded, and computation is wasted on redundant global search. The third is trajectory/noise search, which perturbs the initial noise or explores alternative sampling paths [xu2023restart, ramesh2025test]. Although more fine-grained, it still operates at the full-image level. Searching over all regions can inadvertently disturb areas that are already of high quality, thereby leading to inefficiency and instability. Despite their differences, all three categories share a fundamental limitation: they operate at the full-image level, as shown in Fig. 1. Consequently, they overlook the inherent spatial heterogeneity of image quality and fail to exploit the potential of localized refinement.
This observation naturally motivates an orthogonal direction: localized TTS, where only defective regions are resampled while preserving high-quality content [cao2025temporal], thereby substantially reducing the search space. Yet turning this idea into practice introduces two key challenges. The first is accurate defect localization: since the distribution of artifacts is complex and prompt-dependent, reliably identifying regions that truly require correction is non-trivial [liu2024improved, zhang2023perceptual]. The second is maintaining consistency in local resampling: as the sampling trajectory is globally defined, locally modifying only a subset of regions may introduce incompatibility, leading to semantic drift, stylistic inconsistency, or boundary artifacts that degrade perceptual quality [song2020score, xu2023restart].
In this paper, we propose LoTTS (Prompt-Guided Localized Test-Time Scaling), the first fully training-free localized TTS framework. Specifically, LoTTS consists of two key components: defect localization and consistency maintenance. For defect localization, LoTTS exploits attention signals from diffusion models to construct a quality-aware mask. We contrast semantically equivalent but quality-differentiated prompts (e.g., “a high-quality image of {p}” vs. “a low-quality image of {p}”) and compute the difference between their cross-attention maps, which highlights candidate defective regions [hertz2022prompt, chefer2023attend, chung2024style]. To suppress spurious activations, these difference maps are reweighted with the original prompt attention, filtering out irrelevant background responses. The resulting signals are then propagated via self-attention to enforce spatial coherence, yielding an automated defect mask without reliance on external predictors or manual annotations. For consistency maintenance, LoTTS introduces a localized defect-aware resampling mechanism. After initial sampling, we inject controlled noise into the defective regions of the generated image at an intermediate timestep, followed by localized denoising [meng2021sdedit]. To reconcile local refinements with preserved content, we perform a few global denoising steps, harmonizing style and structure across the image. This ensures that defects are corrected without sacrificing global fidelity. Furthermore, LoTTS is fully plug-and-play and can be seamlessly applied to both diffusion- and flow-based generative models [song2020score, lu2022dpm]. Our main contributions are summarized as follows:
  • We propose LoTTS, the first training-free localized Test-Time Scaling framework, which shifts the focus of inference-time compute from the entire image to defect regions, addressing the inefficiency of global search.
  • We introduce an efficient localized noise search mechanism, comprising prompt-driven quality map generation and localized defect-aware resampling.
[…]
Figure 2: Overview of LoTTS. LoTTS begins by generating multiple candidate images from different noise seeds, identifies defective regions via high-/low-quality prompt contrast, constructs a quality-aware defect mask, injects noise selectively within the masked areas, performs localized denoising with spatial and temporal consistency, and finally selects the best refined sample with a verifier. This pipeline selectively improves low-quality regions while preserving global fidelity.
2 Related work
Image Generation and Evaluation.   Image generation has advanced from GANs [goodfellow2020generative, karras2019style] and autoregressive models [van2016pixel, van2016conditional] to flow-based and diffusion models. GANs achieve high fidelity but are unstable, while autoregressive models capture dependencies yet remain slow. Early flows [kingma2018glow] provide exact likelihoods but scale poorly, whereas recent flow matching and rectified flow methods [lipman2022flow, albergo2023stochastic, kim2025inference] connect closely to diffusion and offer efficient alternatives. Diffusion models [ho2020denoising, rombach2022high] now dominate, delivering state-of-the-art text-to-image synthesis [rombach2022high, podell2023sdxl]. In parallel, evaluation methods have evolved: earlier work often collapsed quality into a single global score [wang2004image, salimans2016improved, heusel2017gans], while spatial indicators relied on texture- or frequency-based heuristics [yu2019attributing, durall2020watch] or classifier- and VLM-based predictors [liu2021aligning, zhang2023perceptual, liu2024improved], typically requiring supervision and external datasets. In contrast, LoTTS leverages the diffusion model’s inherent attention to localize defects automatically, enabling training-free, localized refinement.
Defect Localization.   A substantial body of work aims to localize defects in generated images. Early methods analyze GAN or super-resolution artifacts through spatial or frequency cues [yu2019attributing, durall2020watch, dzanic2020fourier, you2025uncovering]. More recent approaches train supervised models to predict perceptual-error or artifact heatmaps [zhang2023perceptual, li2024g, ren2025otsurv], or pursue defect detection via uncertainty estimation [kou2023bayesdiff] and diffusion-trajectory analysis [cao2025temporal]. In contrast to methods that rely on external predictors, supervised training, or heavy post-hoc analysis, LoTTS derives defect localization directly from the diffusion model itself. Our masks are computed from attention responses to positive and negative prompts, leveraging the model’s inherent cross-modal alignment. This makes LoTTS fully training-free and inherently consistent with the generative semantics of the T2I model.
Generation Quality Enhancement.   Image Editing in modern T2I systems is typically implemented as a post-processing stage. The cascade paradigm, consisting of a generator followed by a refiner, has become a standard design, as in IF [genIF, you2022megan], SDXL [podell2023sdxl] and SD Cascade [pernias2023wurstchen], where an additional pass refines the entire image to enhance global fidelity. However, such refiners usually require separate training and introduce non-negligible computational overhead. In contrast, SDEdit [meng2021sdedit] shows that diffusion models can perform localized edits via noise–denoise updates within user-specified masks, suggesting the potential of training-free localized refinement. This localized resampling mechanism inspires our approach, which extends SDEdit’s manual editing to automated quality-aware refinement at test time.
Test-Time Scaling in Vision.   Some test-time scaling methods have been proposed to enhance diffusion model generation by allocating more computation at inference. Early work simply increased denoising steps, but improvements saturate quickly beyond a certain number of function evaluations (NFE) [karras2022elucidating, song2020denoising, song2020score]. Recent studies therefore explore alternative directions, such as Best-of-$N$ search, where multiple candidates are generated from different noise seeds and a verifier selects the best one [liu2024correcting, wang2023exploring], searching over noise initializations [song2020score, xu2023restart], optimizing sampling trajectories [ramesh2025test] with verifier feedback [song2020denoising, karras2022elucidating, liu2022pseudo, lu2022dpm, salimans2022progressive], or adopting evolutionary [he2025scaling] and tree-search methods [yoon2025fast]. Unlike these approaches, which all operate on the entire image and require full regeneration, our LoTTS performs localized TTS by concentrating on low-quality regions for greater efficiency.
3 Preliminary
Diffusion Models.   Diffusion models transform a simple source distribution, e.g.a standard Gaussian, into a target data distribution $p_{0}$. In diffusion models [sohl2015deep, ho2020denoising], the forward process gradually corrupts clean data with noise, as
$\bm{x}_{t}=\alpha(t)\bm{x}_{0}+\sigma(t)\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (1)
where $\alpha(t)$ and $\sigma(t)$ denote the noise schedule, and $t\in[0,T]$. To recover data from its diffused representation, diffusion models generally rely on Stochastic Differential Equation(SDE)-based sampling during inference [song2020score, song2020denoising], which introduces stochasticity at every denoising step:
$\bm{x}_{t-1}$ $=\sqrt{\alpha(t-1)}\left(\frac{\bm{x}_{t}-\sqrt{1-\alpha(t)}\,\epsilon_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha(t)}}\right)$
$\quad+\sqrt{1-\alpha(t-1)-\sigma(t)^{2}}\,\epsilon_{\theta}(\bm{x}_{t},t)$
$\quad+\sigma(t)\,\epsilon(t).$ (2)
Here, $\epsilon_{\theta}(\bm{x}_{t},t)$ denotes the predicted noise at step $t$, $\alpha(t)$ and $\sigma(t)$ are the noise schedule parameters, and $\epsilon_{t}\sim\mathcal{N}(0,I)$ is a standard Gaussian noise.
Flow Models.   Flow models [lipman2022flow, albergo2023stochastic] parameterize the velocity field $u_{t}\in\mathbb{R}^{d}$ and generate samples by solving the Flow Ordinary Differential Equation (ODE) [song2020score] backward from $t=T$ to $t=0$:
$d\bm{x}_{t}=u_{t}(\bm{x}_{t})\,dt,$ (3)
where $u_{t}(\bm{x}_{t})$ denotes the velocity field learned by the flow model, and $dt$ is an infinitesimal step along the reverse-time ODE. This deterministic dynamics evolves $\bm{x}_{t}$ continuously in time, producing identical outcomes for the same input and limiting the applicability of test-time scaling methods [kim2025inference], which require stochasticity to explore diverse trajectories. To address this limitation, recent studies propose that the deterministic Flow-ODE could be reformulated into an equivalent SDE [albergo2023stochastic, ma2024sit, patel2024exploring, kim2025inference, singh2024stochastic]. The resulting stochastic process can be written as:
$d\bm{x}_{t}=\left(u_{t}(\bm{x}_{t})-\tfrac{\sigma(t)^{2}}{2}\nabla\log p_{t}(\bm{x}_{t})\right)dt+\sigma(t)d\bm{w},$ (4)
where the score function $\nabla\log p_{t}(\bm{x}_{t})$ can be estimated from the velocity field $u_{t}$ [singh2024stochastic]), and the Brownian motion term $d\bm{w}$ introduces stochasticity at each sampling step. This enables LoTTS to naturally apply to Flow-Models.
4 Method
4.1 Preliminaries
Motivation and Challenges.   Existing test-time scaling methods, such as Best-of-$N$ or trajectory search, apply uniform resampling to the entire image, overlooking spatial heterogeneity in quality and thereby wasting computation or even degrading well-formed regions. This observation motivates the consideration of a more targeted form of TTS, one that selectively resamples only defective regions while preserving high-quality areas. By reducing the effective search space and avoiding unnecessary perturbation to clean regions, such a localized TTS paradigm could potentially improve both efficiency and stability compared with full-image resampling. However, realizing this paradigm raises two key challenges: (1) Defect Localization: how to generate reliable resample masks that identify low-quality regions, and (2) Consistency Maintenance: how to perform localized resampling within these regions without disrupting the rest of the image.
Overview of LoTTS.   To address these challenges, we introduce LoTTSLocalized Test-Time Scaling – a fully training-free framework that operationalizes localized resampling for diffusion models. Figure 2 illustrates the LoTTS pipeline. Given a text prompt, LoTTS begins by sampling several noise seeds and generating multiple initial images via standard diffusion denoising. For each initial image, it detects artifact-prone regions using contrastive attention derived from high-quality and low-quality auxiliary prompts, producing a binary mask that highlights potential defects. LoTTS then injects noise exclusively within the masked regions at an intermediate timestep $t_{0}$ and performs localized denoising to refine these areas while keeping the unmasked content unchanged, thereby maintaining global semantic coherence and avoiding boundary artifacts. All refined variants are subsequently evaluated by an external verifier, and the candidate with the highest score is chosen as the final output.
For completeness, pseudocode is given in Appendix B, and Appendix G provides a theoretical analysis showing that LoTTS can be proven to achieve a higher expected quality gain under stated conditions.
4.2 Defect Localization
A key challenge in localized test-time scaling is to identify which regions of a generated image truly require refinement. Since image degradations are usually local and diverse, a reliable defect localization mechanism is indispensable. Existing verifiers typically provide only global quality scores without spatial resolution, making them unsuitable for localized resampling. We leverage the intrinsic attention signals of diffusion models to automatically infer defect regions. The core intuition is that cross-attention with a “low-quality“ prompt points to artifact-prone regions, and comparing it with a “high-quality“ prompt makes those local defects stand out. To obtain stable and semantically meaningful masks, we further propagate and reconcile these information flows.
As illustrated in Fig. 3, the process has four stages: (1) Prompt-driven Discrimination: compute cross-attention maps under high-/low-quality prompts and take their difference to highlight candidate defect regions. (2) Context-aware Propagation: refine these signals by propagating across spatially similar positions, mitigating noise, and enforcing local coherence. (3) Semantic-guided Reweighting: combine with original prompt attention to suppress irrelevant background. (4) Quality-aware Mask Generation: binarize by percentile to control the resampling ratio.
Figure 3: LoTTS Framework for Defect Localization. The pipeline consists of four stages: prompt-driven discrimination, context-aware propagation, semantic-guided reweighting, and quality-aware mask generation.
Prompt-driven Discrimination.   The intuition is that cross-attention maps reflect how different prompts attend to image regions: a “low-quality” prompt tends to focus on artifact-prone areas, while a “high-quality” prompt attends to cleaner regions. By contrasting the two, we can highlight potential defects.
Following the common setting [liu2024towards, helbling2025conceptattention], we obtain a spatial cross-attention vector $\mathrm{CA}\in\mathbb{R}^{S}$ by averaging over tokens, heads, and selected layers, where $S=H_{s}\times W_{s}$ denotes the number of spatial positions. For a given prompt $p$, we construct three variants: a positive prompt (“A high-quality image of {pp}”), a negative prompt (“A low-quality image of {pp}”), and the origin prompt (the original $p$). This yields three attention vectors $\mathrm{CA}_{\mathrm{pos}},\mathrm{CA}_{\mathrm{neg}},\mathrm{CA}_{\mathrm{orig}}\in\mathbb{R}^{S}$. We then define a contrastive cross-attention map as:
$\mathrm{CA}_{\mathrm{diff}}=\mathrm{CA}_{\mathrm{neg}}-\mathrm{CA}_{\mathrm{pos}},$ (5)
which emphasizes spatial locations where the “bad” prompt receives higher attention than the “good” prompt. The origin vector $\mathrm{CA}_{\mathrm{orig}}$ will later serve as a foreground prior in our aggregation step.
Context-aware Propagation.   While the contrastive attention map $\mathrm{CA}_{\mathrm{diff}}$ highlights defect-prone regions, it is often noisy and fragmented, with neighboring pixels showing inconsistent scores. Intuitively, spatially or semantically similar regions should share similar quality signals. To enforce such coherence, we propagate the attention scores using a self-attention matrix derived from the image queries $Q$:
$\mathrm{CA}^{\prime}=\mathrm{SA}\times\mathrm{CA},\qquad\mathrm{SA}=\mathrm{Softmax}{(\frac{QQ^{\top}}{\sqrt{d}})}.$ (6)
This operation smooths the raw attention map by diffusing scores across related spatial positions. Applying it to both $\mathrm{CA}_{\mathrm{diff}}$ and $\mathrm{CA}_{\mathrm{orig}}$ yields refined maps $\mathrm{CA}^{\prime}_{\mathrm{diff}}$ and $\mathrm{CA}^{\prime}_{\mathrm{orig}}$ that are more stable and spatially coherent.
Semantic-guided Reweighting.   Although the contrastive map $\mathrm{CA}^{\prime}_{\mathrm{diff}}$ can reveal defect-prone areas, it often assigns high scores to background regions with little semantic content (e.g., large sky areas). The intuition is that truly meaningful defects should also lie within the semantic foreground defined by the original prompt. To encode this prior, we combine the contrastive map with the original attention map $\mathrm{CA}^{\prime}_{\mathrm{orig}}$, which serves as a soft foreground mask:
$P=\mathrm{CA}^{\prime}_{\mathrm{diff}}+\lambda\,\mathrm{CA}^{\prime}_{\mathrm{orig}},$ (7)
where $\lambda$ balances the defect signal and the foreground prior. The resulting map $P$ emphasizes semantically relevant regions, leading to more reliable defect localization.
Quality-Aware Mask Generation.   Finally, we need to convert the aggregated quality map into a binary mask that specifies which regions should be resampled. The key idea is to only refine the most degraded areas rather than the entire image. To achieve this, we keep the top $r$ proportion of spatial positions with the highest defect scores, ensuring that resampling is both targeted and controllable. Formally, the mask $M\in\{0,1\}^{S}$ is given by:
$M=\mathbb{I}\!\left(P>\mathrm{Perc}(P,\,1-r)\right),$ (8)
where $\mathrm{Perc}(P,\,1-r)$ is the (1r)(1-r)-quantile of the quality map $P$, and $r\in(0,1)$ controls the fraction of resampled regions. The mask is then reshaped into the spatial grid $\mathbf{M}\in\{0,1\}^{H_{s}\times W_{s}}$ to guide localized refinement. We compute the attention mask at $t=0$, matching the final image.
Table 1: Quantitative results on three benchmarks (Pick-a-Pic, DrawBench, COCO2014). Across human-aligned metrics (HPS, AES, Pick, IR) and automated metrics (FID, CLIP), LoTTS consistently outperforms Resampling, Best-of-$N$, and Particle Sampling baselines.
Model Method Pick-a-Pic DrawBench COCO2014
[…]
SD2.1 Resampling 20.44 5.377 20.32 0.236 21.34 5.456 20.23 0.244 15.33 0.201
Best-of-NN 21.56 5.534 21.04 0.470 22.45 5.589 20.59 0.446 13.21 0.252
Particle Sampling 23.44 5.980 21.30 0.530 22.19 5.790 21.23 0.520 12.34 0.260
LoTTS (Ours) 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698 10.89 0.263
SDXL Resampling 23.44 6.011 21.18 0.680 23.84 6.034 21.09 0.657 9.56 0.234
Best-of-NN 24.54 6.198 22.01 0.790 25.27 6.238 22.23 0.756 8.34 0.268
Particle Sampling 25.33 6.235 22.05 0.865 26.46 6.233 22.31 0.844 7.99 0.271
LoTTS (Ours) 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111 7.33 0.297
FLUX Resampling 29.34 6.298 22.07 1.038 29.28 6.223 22.05 1.100 7.01 0.282
Best-of-NN 30.23 6.299 22.89 1.235 30.46 6.290 22.33 1.221 6.34 0.306
Particle Sampling 31.56 6.532 23.31 1.450 32.28 6.523 22.90 1.445 6.02 0.332
LoTTS (Ours) 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623 5.31 0.351
4.3 Consistency Maintenance
With reliable defect masks from the previous step, the next challenge is how to resample the identified regions without disrupting the rest of the image. Restricting updates only to masked regions often introduces boundary artifacts or semantic drift. To address these issues, LoTTS maintains both spatial and temporal consistency during refinement.
Spatial Consistency.   Resampling in diffusion models typically begins by perturbing the latent representation. If perturbation is applied only within the mask, the noise distribution becomes inconsistent with surrounding regions, creating visible seams. We avoid this by injecting comparable noise into both masked and unmasked areas, which balances noise levels and ensures smooth transitions across boundaries. Formally, given the clean latent $\mathbf{x}_{0}$ and binary mask $\mathbf{M}$, we initialize the perturbed latent at timestep $t_{0}$ as:
$\mathbf{x}_{t_{0}}=\alpha(t_{0})\mathbf{x}_{\mathbf{0}}+\sigma(t_{0})((1-\mathbf{M})\odot\mathbf{z}_{\mathbf{bg}}+\mathbf{M}\odot\mathbf{z}_{\mathbf{mask}})$ (9)
where $\mathbf{z}_{\mathbf{bg}}\!\sim\!\mathcal{N}(0,I)$ and $\mathbf{z}_{\mathbf{mask}}\!\sim\!\mathcal{N}(0,I)$.
Temporal Consistency.   A second challenge is preserving the global semantics of the image. Restarting from pure noise discards structure and forces the model to regenerate the entire scene. Instead, following meng2021sdedit, we resume denoising from an intermediate step $t_{0}$, so that global content is retained while localized corrections remain possible. The masked reverse update is:
$\begin{split}\mathbf{x}_{t-\Delta t}=&\,(1-\mathbf{M})\odot\big(\alpha(t)\mathbf{x}_{0}+\sigma(t)\mathbf{z}\big)\\ &+\mathbf{M}\odot\Big(\mathbf{x}_{t}-\epsilon^{2}\epsilon_{\theta}(\mathbf{x}_{t},t)+\epsilon\mathbf{z}\Big).\end{split}$ (10)
where $\epsilon_{\theta}(\mathbf{x}_{t},t)$ denotes the predicted noise, $\epsilon$ is the per-step noise scale, $N$ is the total number of reverse steps, $\sigma(t)$ is the variance schedule at step $t$, $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\Delta t=t_{0}/N$. This keeps unmasked regions faithful while allowing masked regions to be selectively refined.
Holistic Consistency.   Even after localized refinement, residual inconsistencies may remain around boundaries. To restore smooth transitions, we apply a brief global denoising sweep. This final step enforces full-image coherence and yields consistent, high-fidelity outputs.
5 Experiments and Analysis
Benchmarks.   We evaluate LoTTS on three benchmarks: Pick-a-Pic [pick], DrawBench [saharia2022photorealistic], and COCO2014 [lin2014microsoft]. Performance is measured by four human-aligned metrics, HPSv2 [wu2023human], PickScore [pick], ImageReward (IR) [ir] (also serving as the verifier), and Aesthetic Score (AES) [aes_laion], together with FID [heusel2017gans] and CLIP Score [radford2021learning] on COCO.
Baselines.  Experiments are conducted on SD2.1 [rombach2022high], SDXL [podell2023sdxl], and Flux.1-schnell [flux2024]. SD2.1 and SDXL are diffusion-based, while Flux.1-schnell is a rectified flow-based model. We compare LoTTS with representative sampling and search methods under matched NFE budgets, including vanilla Resampling, Best-of-$N$, and Particle Sampling [kim2025test, singhal2025general].
Implementation Details.   Unless otherwise specified, LoTTS uses a mask ratio of $r=50\%$, an attention reweighting coefficient $\lambda=0.5$, a sampler-dependent noise-injection step $t_{0}$ (25/15/1 for SD2.1/SDXL/FLUX), and a default sample count of $N=9$ per prompt, obtained from $s=3$ initial images and $k=2$ localized refinement candidates for each. More implementation details are provided in Appendix A.
Figure 4: Qualitative results on challenging text-to-image prompts. Compared with Resampling, Best-of-NN and Particle Sampling, LoTTS better follows complex prompts. The green borders indicate high-quality generations, red mark lower-quality ones, and brown shows the initial image used as the starting point of localized refinement.
5.1 Main Results
Quantitative Performance.   Tab.1 shows that LoTTS consistently outperforms Resampling, Best-of-$N$, and Particle Sampling[kim2025test, singhal2025general] under matched NFE budgets across SD2.1, SDXL, and FLUX, achieving clear improvements on both human-aligned metrics (HPS, AES, Pick, IR) and automated metrics. While some baselines occasionally obtain slightly higher AES due to the metric’s preference for stylized, high-contrast renderings, LoTTS focuses on structural and semantic correctness, making small AES fluctuations expected. Importantly, the three benchmarks vary widely in dataset scale and prompt distribution, indicating that LoTTS generalizes well across different evaluation settings. Moreover, the results show that LoTTS is architecture- and sampler-agnostic: it works well with both U-Net-based and Transformer-based diffusion model and suitable for both SDE- and ODE-based samplers.
Qualitative Comparison.   Figure 4 shows that LoTTS produces cleaner and more coherent images across challenging prompts. The w/o Scaling results often contain severe artifacts. For example, muddy textures, incomplete patterns, distorted metal parts, and missing wolf for those prompts. Best-of-$N$ and Particle Sampling substantially improve quality but still suffer from issues such as inconsistent lighting, oversaturated colors, and local geometric distortions. In contrast, LoTTS performs targeted local refinements on the initial generation (with brown border), improving semantic details while preserving the global layout. These observations demonstrate that LoTTS more effectively improves visual fidelity and stability than existing TTS baselines.
Figure 5: Localized Refinement Results on SD2.1. LoTTS corrects diverse localized artifacts, such as distorted object geometry, duplicated fingers, texture noise, and warped text. Circles highlight corrected regions.
Local Refinement Capability.   Fig. 5 illustrates the different types of localized artifacts that LoTTS effectively corrects. The cow example shows a distorted or missing head; the hand example exhibits duplicated or fused fingers; the bicycle contains wavy, noisy textures on the spokes; and the airplane displays warped, unreadable text on the fuselage. Across these diverse cases, LoTTS reliably refines the corrupted regions while preserving surrounding content, demonstrating both the effectiveness of localized resampling and its applicability to a broad range of artifact types. White and Red circles both mark the corrected areas.
More Results.   More qualitative results and failure cases provided in Appendix C and F.
5.2 Ablations
Mask Generation.   Fig.6 visualizes the intermediate outputs of our mask generation pipeline, showing how each stage contributes to coherent defect localization. Remarkably, despite being unsupervised, our final mask qualitatively appears more precise and semantically aligned than the supervised defect-detection model [li2024g]. As summarized in Fig.7, removing any component – whether the discriminative difference map $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, the propagated original attention $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, or the context-aware propagation – consistently degrades all four metrics. The variant without $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ performs the worst overall, indicating the importance of the strong semantic foreground prior. The no-mask and random-mask baselines also underperform, further confirming the need for accurate, structure-aware masks.
Figure 6: Visualization of Cross-Attention–based Defect Localization. From the original ($\mathrm{CA}_{\mathrm{orig}}$), positive ($\mathrm{CA}_{\mathrm{pos}}$), and negative ($\mathrm{CA}_{\mathrm{neg}}$) prompts, LoTTS derives discriminative differences ($\mathrm{CA}_{\mathrm{diff}}^{\prime}$) and combines them with propagated original attention ($\mathrm{CA}_{\mathrm{orig}}^{\prime}$) to produce a coherent defect mask ($\mathrm{CA}_{\mathrm{diff}}^{\prime}+\lambda\mathrm{CA}_{\mathrm{orig}}^{\prime}$) that highlights low-quality regions. The Reference shows the masks from a suboptimal supervised model [li2024g].
Figure 7: Ablation of Mask Generation for LoTTS. Mask generation ablations show each component contributes to generation performance.
Quality Prompt Design.   As shown in Tab. 2, different auxiliary prompts (e.g., “a good/bad image of {p}”) yield similar results, indicating that LoTTS is robust to prompt phrasing. These prompts serve only to elicit contrastive attention patterns – “good” prompts highlight regions the model considers well-formed, while “bad” prompts emphasize areas linked to low quality. The consistent performance across phrasing suggests that LoTTS relies on this quality contrast rather than the specific wording, and that diffusion models encode a stable notion of visual quality that our LoTTS can reliably exploit.
Table 2: Ablation of Prompt Design for LoTTS. “Good/Bad” uses the prompts “a good/bad image of {p}”. “High/Low-quality” uses “a high/low-quality image of {p}”.
[…]
Good/Bad 6.29 28.62 22.21 1.140
High/Low-quality 6.29 28.69 22.29 1.137
Figure 8: Parameter sensitivity analysis of LoTTS. Varying localized refinement iterations ($k$), resample area ratio ($r$), and denoising start step ($t_{0}$) shows that LoTTS maintains stable improvements across HPS, AES, Pick, and IR metrics.
Number of Refinements ($k$).   As shown in Fig. 8, performance peaks at $k=2$. Due to the stochastic nature of diffusion sampling, a single refinement may miss the correct reconstruction, while $2$ attempts provide enough diversity to reliably fix most artifacts. Beyond this point, additional refinements rarely offer new solutions, suggesting that the effective correction space is reached within just a few attempts.
Mask Area Ratio ($r$).   Moderate mask ratio r=50r=50% works best. Small masks tend to miss parts of the defects, while large masks begin to overwrite clean regions and behave like full-image resampling. Overall, LoTTS is fairly robust to the choice of $r$, showing stable performance across a wide range of mask sizes.
Noise Injection Step ($t_{0}$).   The optimal noise injection step depends on the model’s generative ability. Injecting noise too early disrupts too much original information, while injecting too late provides too little flexibility to correct defects, so $t_{0}$ must balance structural preservation with refinability. For diffusion models like SD2.1 and SDXL, mid-range steps work well, whereas FLUX converges very early and therefore benefits from injecting noise closer to the noise-dominated steps.
Scaling Efficiency.   As shown in Fig. 9, LoTTS reaches Best-of-$N$ performance with far fewer samples, yielding $2.8\times$$3.3\times$ speedups on SD2.1 and SDXL and up to $4\times$ on FLUX. The larger speedup on stronger models shows that even high-quality generators still leave local defects that LoTTS can efficiently fix, positioning localized TTS as an increasingly essential tool in the next generation of large generative models.
Figure 9: Scaling Comparison on LoTTS vs. Best-of-NN Sampling.
More Ablations.   Additional ablation and parameter sensitivity results are included in Appendix D and E.
5.3 Discussion
As shown in Fig.10, LoTTS occasionally produces failures. We think LoTTS failures mainly fall into three categories. First, imperfect masks: extremely subtle defects or high-frequency regions may cause missed or unstable attention signals. This challenge is not unique to our training-free LoTTS – supervised defect-detection models exhibit similar limitations – but importantly, our Appendix G shows that localized TTS can outperform global resampling even with imperfect masks. Second, global defects: issues involving perspective, spatial relations, or global semantics cannot be fixed through local resampling alone, pointing toward future work on adaptive local–global TTS strategies. Third, initial images at quality extremes: low-quality images may be too fragile to refine locally, while near-perfect images offer little room for improvement, suggesting that early-stopping mechanisms may further improve LoTTS. Overall, these cases point to natural opportunities for extending localized TTS, and LoTTS marks an early step toward more adaptive scaling methods.
Figure 10: Failure cases of LoTTS. Left: local refinement introduces a facial distortion in the cat image. Right: the image has minimal visible change after refinement.
6 Conclusion
We proposed LoTTS, a training-free framework that extends test-time scaling from global resampling to localized refinement. Unlike conventional methods that apply sampling uniformly, LoTTS leverages defect-aware masks and consistency constraints to focus computation where it matters most. This design not only improves image quality and efficiency but also shows that scaling can be made adaptive to the spatial heterogeneity of generative outputs. Beyond diffusion models, the principle of region-aware scaling provides a general perspective for developing more efficient and controllable inference strategies in generative AI. Our analysis further establishes theoretical conditions under which localized scaling provably outperforms global sampling, and we believe LoTTS opens up promising directions for integrating fine-grained control into test-time algorithms across diverse generative architectures.
[…]
Supplementary Material
AImplementation Details...........................................................................................................A
BAlgorithmic Details...........................................................................................................B
  B.1Notation and Preliminaries...........................................................................................................B.1
  B.2Defect Localization Procedure...........................................................................................................B.2
  B.3Localized Resampling Procedure...........................................................................................................B.3
  B.4Overall Algorithm of LoTTS...........................................................................................................B.4
CQualitative Results...........................................................................................................C
  C.1Mask Generation Examples...........................................................................................................C.1
  C.2Localized Refinement Examples...........................................................................................................C.2
DAblations...........................................................................................................D
  D.1Mask Generation...........................................................................................................D.1
  D.2Prompt Design...........................................................................................................D.2
EParameter Analysis...........................................................................................................E
  E.1Hyperparameter Sensitivity...........................................................................................................E.1
  E.2Scaling Comparison with Best-of-$N$...........................................................................................................E.2
FDiscussion...........................................................................................................F
  F.1Failure Analysis...........................................................................................................F.1
  F.2Future Directions...........................................................................................................F.2
GTheoretical Analysis...........................................................................................................G
  G.1Problem Setup and Preliminaries...........................................................................................................G.1
  G.2Main Theorem...........................................................................................................G.2
  G.3Derived Corollaries...........................................................................................................G.3
  G.4Practical Q&A...........................................................................................................G.4
Appendix A Implementation Details
Hardware and software.   All experiments were run on NVIDIA A800 (80GB) GPUs; most runs use a single GPU unless otherwise stated. Mixed precision is enabled via torch.cuda.amp.autocast (fp16) unless the metric implementation requires fp32.
Checkpoints and preprocessing.   We use the official public checkpoints for SD2.1, SDXL, and FLUX.1-schnell. Images are generated at $512{\times}512$ for SD2.1 and at $1024{\times}1024$ for SDXL and FLUX. For latent diffusion backbones, we use the default VAE shipped with each checkpoint. Evaluation images are fed to each metric with the metric’s own preprocessing. Unless noted, we do not run SDXL-Refiner to avoid confounding the effect of local resampling.
Default sampler and schedule.   For SD2.1, we use SDE-DPM-Solver++ sampling with 50 NFEs, and CFG scale =7.5=7.5. For SDXL, we use SDE-DPM-Solver++ sampling with 30 NFEs, and CFG scale =5.5=$5.$ For FLUX, we use modified SDE-DPM-Solver++ sampling with 10 NFEs, and no CFG. For LoTTS, we inject noise to reach an intermediate step $t_{0}$ (Sec. 4), then perform local masked refinement followed by a short global integration phase. We keep the original prompt and CFG weight unchanged during refinement.
Attention hook and mask resolution.   For SD2.1, we hook U-Net cross-attention tensors at the $16{\times}16$ spatial resolution blocks (for $512{\times}512$ images). For SDXL, we hook U-Net cross-attention tensors at the $64{\times}64$ spatial resolution blocks (for $1024{\times}1024$ images). For FLUX, we extract Transformer cross-attention tensors from the $64{\times}64$ resolution of the last 10 blocks (for $1024{\times}1024$ images), and generate attention maps following the method of [helbling2025conceptattention]. The attention maps are averaged across heads, mean-pooled across the feature dimension, and then averaged across tokens to obtain a single spatial map per prompt type (pos/neg/origin). Self-attention maps are normalized to $[0,1]$. The quality map $P$ uses equation 7 in the paper with $\lambda=0.5$ by default. We compute attention and the final mask at t=0t{=}0 (last step), where maps and image latents are already stable. Masks are kept in latent resolution during resampling; only for visualization we bilinearly upsample to pixel space.
Datasets.   For Pick-a-Pic, we follow the official prompt pool for preference-oriented evaluation. For DrawBench, we use all categories and report aggregate metrics. For COCO2014, we use the widely-used COCO2014-30k.
Metrics.   We report HPSv2, PickScore, ImageReward, and AES using the authors’ official weights and preprocessing. CLIP score and FID are computed with standard settings: CLIP uses ViT-L/14 unless otherwise stated; FID uses Inception-V3 features, $2048$-D activations. ImageReward (IR) score is used as the global selection score for Best-of-$N$, LoTTS, and other search-based baselines.
Baselines and compute budgets.   We compare with Resample, Best-of-N and Particle Sampling under matched NFEs. For “Best-of-NN” baselines, we sample $N$ candidates from scratch and select by the same metric used for reporting to avoid selection bias. For LoTTS we use $k=2$ localized refinement iterations by default.
Reproduction checklist.   (1) Use the same checkpoint and VAE; (2) match resolution, sampler, steps, CFG; (3) compute the mask at t=0t{=}0 with the same prompt templates; (4) keep $\lambda$, $r$, $t_{0}$, $T^{\prime}$, $k$ as in Tab. 1; (5) fix seeds and determinism flags; (6) run the same metric code and preprocessing.
SD 2.1 SDXL FLUX
Res. $512{\times}512$ ${\times}1024 1024×1024$${\times}1024$
Sampler SDE-DPM- Solver++ SDE-DPM- Solver++ Modified SDE- DPM-Solver++
Steps 50 30 10
CFG 7.5 5.5 None
$\lambda$ 0.50 0.50 0.50
$t_{0}$ 25 15 1
$T^{\prime}$ 50 30 10
[…]
Table 1: Default LoTTS hyperparameters. Transposed single-column version with wrapped sampler names.
Appendix B Algorithmic Details
This section presents the full algorithmic components of LoTTS. Algorithm 1 extracts a binary resample mask $M$ from prompt-conditioned cross-attention maps. Algorithm 2 integrates this mask into a localized resampling module that refines only defect-prone regions while preserving background structure. Algorithm 3 combines these components into a depth-first search (DFS) that explores global diversity and localized refinement, ultimately selecting the highest-scoring image.
B.1 Notation and Preliminaries
Tab. 2 summarizes all symbols used in Algorithms 1, 2, and 3. The goal is to unify notation so that every variable appearing in the pseudocode is well-defined. We explicitly distinguish quantities related to user prompts, mask generation, localized refinement, and DFS-level search dynamics.
Symbol Description
$p$ original user prompt
$p_{\text{pos}},p_{\text{neg}}$ quality-aware prompts (“a high-quality image of $\{p\}$”, “a low-quality image of $\{p\}$”)
$x$ current image passed through DFS recursion
$x_{\mathrm{in}},x_{\mathrm{out}}$ input/output image of LocalizedResample
$v_{\mathrm{out}}$ verifier score of $x_{\mathrm{out}}$
$x_{\text{best}},v_{\text{best}}$ current best image and score during DFS search
$x^{\star}$ final best refined image returned by DFS
$M$ binary resample mask from Algorithm 1 (latent resolution)
$r$ area ratio used in quantile thresholding
$\mathcal{L}$ set of U-Net layers used to extract cross-attention maps
$\lambda$ foreground weighting coefficient in mask reweighting
$T$ total number of diffusion steps (scheduler alignment)
$t_{0}$ intermediate re-noise step for localized resampling
$t_{g}$ global integration step
$S$ number of global seeds at DFS depth 1
$K$ number of localized refinements per global seed
$D$ DFS maximum depth (D=2D{=}2)
NN number of refinement steps (also for $t_{g}\!\to\!0$ discretization)
$\alpha(t),\sigma(t)$ noise schedule for both forward and reverse updates
$\epsilon_{\theta}(\cdot)$ noise predictor (with CFG if enabled)
$V(\cdot)$ global verifier producing a scalar score
SampleBase base sampler generating an initial global sample conditioned on $p$
$\mathrm{CrossAttn}$ cross-attention maps from U-Net layers $\mathcal{L}$ during a forward pass with latent $x$ and prompt $p$
Table 2: Unified notation for Algorithms 1, 2, and 3.
B.2 Defect Localization Procedure
Algorithm 1 details the MaskGen module, whose goal is to identify regions likely to contain visual defects. The key intuition is that low-quality prompts activate areas where the model is susceptible to producing artifacts, whereas high-quality prompts suppress them. By contrasting their cross-attention maps and propagating information through self-attention layers, MaskGen amplifies quality-sensitive regions while preserving semantic coherence. Quantile thresholding is then applied to select only the top-$r$ most suspicious patches. The resulting binary mask $M$ provides spatial guidance for localized refinement.
Algorithm 1 MaskGen: Defect Localization
1:Current image $x$, prompt $p$, positive prompt $p_{\text{pos}}$, negative prompt $p_{\text{neg}}$, selected layers $\mathcal{L}$, weight $\lambda$, area ratio $r$
2:Binary mask $M$
3:Prompt-driven Discrimination
[…]
8:Context-aware Propagation
[…]
11:Semantic-guided Reweighting
[…]
13:Mask Generation
[…]
15:return $M$
B.3 Localized Resampling Procedure
Algorithm 2 defines LocalizedResample, which refines only the defective regions indicated by mask $M$. The process begins by re-noising the entire image to an intermediate timestep $t_{0}$, ensuring that both masked and unmasked regions lie on the diffusion trajectory. During the refinement phase ($t_{0}\!\rightarrow\!t_{g}$), masked regions perform reverse denoising steps guided by $\epsilon_{\theta}$, while unmasked regions follow a scheduled forward-noise injection that preserves background content. A final integration phase ($t_{g}\!\rightarrow\!0$) restores global coherence using standard reverse diffusion. The verifier $V$ evaluates the refined output, providing a scalar score used in DFS.
Algorithm 2 LocalizedResample: Attention-Guided Resampling (single refinement)
1:Input image $x_{\mathrm{in}}$ (at t=0t{=}0), prompt $p$, total steps $T$, re-noise step $t_{0}$, integration step $t_{g}$, mask $M$, diffusion model, verifier $V$
2:Refined image $x_{\mathrm{out}}$, score $v_{\mathrm{out}}$
[…]
5:Global re-noise to $t_{0}$
6:Sample $z_{\text{bg}},z_{\text{mask}}\sim\mathcal{N}(0,I)$
[…]
8:Masked refinement ($t_{0}\rightarrow t_{g}$)
9:for $t=t_{0},\,t_{0}-\Delta t,\,\dots,\,t_{g}$ do
10:   Sample $z_{t}\sim\mathcal{N}(0,I)$
11:   $x_{t-\Delta t}\leftarrow(1-M)\odot(\alpha(t)x_{\text{anchor}}+\sigma(t)z_{t})+M\odot(x_{t}-\Delta t\,\epsilon_{\theta}(x_{t},t;p)+\sigma(t)z_{t})$
12:end for
13:Global integration ($t_{g}\rightarrow 0$)
14:for $t=t_{g},\,t_{g}-\Delta t,\,\dots,\,0$ do
15:   Sample $z_{t}\sim\mathcal{N}(0,I)$
16:   $x_{t-\Delta t}\leftarrow x_{t}-\Delta t\,\epsilon_{\theta}(x_{t},t;p)+\sigma(t)z_{t}$
17:end for
[…]
20:return $x_{\mathrm{out}},\,v_{\mathrm{out}}$
21:$T$ is included for scheduler compatibility; discretization uses $\Delta t$.
B.4 Overall Algorithm of LoTTS
Algorithm 3 presents the full LoTTS pipeline expressed as a depth-first search with depth D=2D{=}2. Depth 0 explores global diversity by sampling $S$ initial images from scratch via SampleBase. At depth 1, each global sample undergoes $K$ localized refinements guided by Algorithm 2. Every candidate is scored by the verifier $V$, and DFS continually tracks the highest score seen so far. This hierarchical search effectively combines global variability with targeted local improvements, achieving strong test-time scaling.
Algorithm 3 Overall Algorithm of LoTTS (DFS with depth D=2D{=}2)
1:Prompt $p$, DFS depth D=2D{=}2, global seeds $S$, localized refinements $K$, total steps $T$, re-noise $t_{0}$, integration $t_{g}$, verifier $V$
2:Best refined image $x^{\star}$
3:function DFS($p,d,x,x_{\text{best}},v_{\text{best}}$)
4:   if $x\neq\varnothing$ then
5:     $v\leftarrow V(x)$
6:     if $v>v_{\text{best}}$ then
7:        $x_{\text{best}}\leftarrow x$
8:        $v_{\text{best}}\leftarrow v$
9:     end if
10:   end if
11:   if $d=D$ then
12:     return $(x_{\text{best}},v_{\text{best}})$
13:   end if
14:   if $d=0$ then // root $\rightarrow$ global seeds
15:     for $i=1$ to $S$ do
16:        $x_{\mathrm{in}}\leftarrow\text{SampleBase}(p)$
17:        $(x_{\text{best}},v_{\text{best}})\leftarrow\textsc{DFS}(p,1,x_{\mathrm{in}},x_{\text{best}},v_{\text{best}})$
18:     end for
19:     return $(x_{\text{best}},v_{\text{best}})$
20:   else// $d=1$: global sample $\rightarrow K$ localized refinements
21:     $M\leftarrow\text{MaskGen}(x,p)$
22:     for $k=1$ to $K$ do
23:        $(x_{\mathrm{out}},v_{\mathrm{out}})\leftarrow\text{LocalizedResample}(x,p,T,t_{0},t_{g},M,V)$
24:        $(x_{\text{best}},v_{\text{best}})\leftarrow\textsc{DFS}(p,2,x_{\mathrm{out}},x_{\text{best}},v_{\text{best}})$
25:     end for
26:     return $(x_{\text{best}},v_{\text{best}})$
27:   end if
28:end function
[…]
30:return $x^{\star}$
Appendix C Qualitative Results
We provide additional qualitative comparisons to highlight LoTTS’s local refinement capabilities.
C.1 Mask Generation Examples
Fig. 1 and Fig. 2 provide a detailed visualization of the mask generation process in SD2.1 and FLUX. Across both models, LoTTS exhibits several consistent behaviors that highlight its ability to identify and isolate low-quality regions.
First, the original cross-attention map ($\mathrm{CA}_{\mathrm{orig}}$) captures semantic foreground structure but provides limited information about actual degradations. The positive and negative prompt attention maps ($\mathrm{CA}_{\mathrm{pos}}$, $\mathrm{CA}_{\mathrm{neg}}$) reveal complementary patterns: $\mathrm{CA}_{\mathrm{pos}}$ tends to emphasize well-formed regions aligned with the prompt, while $\mathrm{CA}_{\mathrm{neg}}$ highlights visually inconsistent or low-confidence areas. Taking their contrast produces $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, which reliably exposes localized artifacts such as missing limbs, distorted geometry, texture noise, or incorrect shading.
Next, the propagated attention $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ spreads semantic information through self-attention layers, sharpening object boundaries and reinforcing regions that are coherent with the global context. When combined ($\mathrm{CA}_{\mathrm{diff}}^{\prime}+\mathrm{CA}_{\mathrm{orig}}^{\prime}$), the resulting mask becomes both defect-aware and semantically grounded: irrelevant background areas are suppressed, while problematic regions remain dominant.
As shown in Fig. 1, SD2.1 masks tend to be smoother and highlight larger defective zones. In contrast, FLUX (Fig. 2) produces finer-grained and more spatially precise masks, reflecting its higher visual capacity and ability to reveal subtle local degradations (e.g., thin structure distortions, glass reflections, or geometric irregularities). Across diverse prompts (human bodies, animals, food, vehicles, and indoor scenes) LoTTS consistently localizes error-prone regions, demonstrating its robustness.
Figure 1: Additional examples of mask generation in SD2.1. LoTTS produces reliable quality-aware masks across diverse prompts, effectively localizing low-quality regions for refinement.
Figure 2: Additional examples of mask generation in FLUX. LoTTS produces finer and more detailed masks in FLUX, consistently localizing local degradations under varied prompts and enabling more precise localized resampling.
C.2 Localized Refinement Examples
Fig. 3 presents additional before–after examples from FLUX, illustrating how LoTTS performs targeted corrections on locally degraded regions while preserving the global layout and scene semantics. In the upper-left example, a tree in the background appears blurry and lacks structure; after refinement, LoTTS restores clear branches and edges without altering the surrounding scene. In the upper-right example, the original image contains an animal with an implausible and distorted body shape. LoTTS corrects the malformed geometry, recovering a coherent and recognizable giraffe with consistent proportions. The lower-left example highlights LoTTS’s ability to enhance fine textures. The bear statue originally exhibits smeared and noisy patterns on its torso. After refinement, LoTTS produces sharper textures and more aesthetically coherent shading, improving the realism of the material. Finally, in the lower-right example, the bowl shows washed-out colors and softened boundaries. LoTTS sharpens the rim, restores the color transitions, and produces a cleaner overall appearance while keeping the composition unchanged. Across all cases, the refined results maintain the original structure and content, modifying only the regions that exhibit degradation. The red and green circles in Fig. 3 indicate the specific areas improved by LoTTS, demonstrating its ability to perform precise, localized enhancements without introducing unintended global changes.
Figure 3: Localized Refinement Examples in FLUX. LoTTS achieves consistent local refinements across diverse prompts, enhancing structural integrity, semantics, and perceptual quality. The red and green circles highlight corrected regions.
Appendix D Ablations
D.1 Mask Generation
We conduct detailed ablations to understand the role of each component in our mask-generation pipeline. Tab. 3 and Fig. 4 summarize results on SD2.1, SDXL, and FLUX. Across all settings, modifying or removing any component degrades performance, but the severity of degradation follows a clear pattern that reflects how each factor affects mask quality and refinement behavior.
Worst case: removing $\mathrm{CA}_{\mathrm{orig}}^{\prime}$.   The largest drop occurs when $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ is removed. Without the original cross-attention prior, the mask loses its semantic foreground constraint and often activates on background regions rather than meaningful objects. This severely reduces mask precision and leads the model to refine irrelevant areas, producing the lowest scores across nearly all metrics. This confirms that $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ is essential for locating semantically valid regions where refinement should be applied.
Random mask: too fragmented and semantically inconsistent.   The second-worst performance comes from using a random mask. Unlike the “w/o Mask” baseline, the random mask still enforces locality, but it does so in an unstructured and highly fragmented manner. Because foreground regions and background pixels are randomly mixed, the mask fails to form coherent spatial blocks, making refinement unstable and preventing the model from effectively correcting artifacts. This produces masks with both low precision and low recall, explaining why the random baseline consistently underperforms.
No mask: global resampling is unnecessarily destructive.   Removing the mask (“w/o Mask”) yields the next level of degradation. In this case, refinement becomes global: the model resamples the entire image, including regions that were already high quality. Although this avoids the fragmentation issue of random masks, the lack of locality causes semantic drift and over-editing, which lowers HPS, AES, and Pick metrics. This demonstrates that locality is crucial for preserving global semantics while improving specific defective regions.
Removing $\mathrm{CA}_{\mathrm{diff}}^{\prime}$: no defect-specific cue.   Mask quality improves when keeping $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ while removing $\mathrm{CA}_{\mathrm{diff}}^{\prime}$. However, without the differential signal, the mask becomes insensitive to quality differences and focuses only on semantic saliency. As a result, some artifacts are missed, lowering recall and reducing the improvement rate. This explains why performance is better than random or no-mask baselines, but still clearly below our full design.
Removing Propagation: masks lack spatial continuity.   The best-performing ablation variant is “w/o Propagation.” The mask still identifies the correct semantic and defective regions thanks to both cross-attention terms, but lacks spatial smoothness. Consequently, the mask is less contiguous and refines regions in a slightly noisy manner. This produces moderate performance drops, indicating that while Propagation is not as critical as the attention signals, it contributes substantially to stable, coherent refinement.
Summary.   Across all variants, the degradation trend follows:
$\text{w/o }\mathrm{CA}_{\mathrm{orig}}^{\prime}$ $<\text{Random Mask}<\text{w/o Mask}$
$<\text{w/o }\mathrm{CA}_{\mathrm{diff}}^{\prime}<\text{w/o Propagation}<\text{Ours}.$
This ordering reflects the relative contributions of semantic grounding from $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, defect sensitivity from $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, and spatial coherence from Propagation. Together, these factors enable LoTTS to produce precise, contiguous, and defect-aware masks that yield the strongest overall refinement performance.
Table 3: Ablation study of mask generation strategies on Pick-a-Pic and DrawBench. Results for SD2.1, SDXL, and FLUX show that removing or altering components (e.g., w/o Mask, Random Mask, w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, w/o $\mathrm{CA}_{\mathrm{diff}}$, or w/o Propagation) degrades performance, confirming the effectiveness of our full design.
Model Mask Strategy Pick-a-Pic DrawBench
[…]
SD2.1 w/o Mask 21.21 5.245 21.14 0.452 22.33 5.591 20.66 0.453
Random Mask 20.59 5.325 21.01 0.457 22.34 5.580 20.67 0.452
w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ 20.43 5.223 20.87 0.444 21.22 5.570 20.44 0.445
w/o $\mathrm{CA}_{\mathrm{diff}}^{\prime}$ 21.31 5.303 20.95 0.458 22.53 5.605 21.01 0.500
w/o Propagation 23.51 5.705 21.12 0.600 22.89 5.811 21.36 0.688
Ours 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
SDXL w/o Mask 24.20 6.120 22.04 0.780 25.31 6.244 22.31 0.788
Random Mask 24.19 6.115 22.01 0.755 25.65 6.231 22.32 0.778
w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ 24.00 6.100 22.00 0.748 24.55 6.222 22.28 0.766
w/o $\mathrm{CA}_{\mathrm{diff}}^{\prime}$ 26.00 6.201 22.05 0.850 26.35 6.272 22.30 0.901
w/o Propagation 28.11 6.290 22.11 1.001 28.30 6.301 22.35 1.100
Ours 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
FLUX w/o Mask 30.25 6.324 22.67 1.241 31.32 6.231 22.20 1.357
Random Mask 30.15 6.313 22.55 1.211 31.23 6.225 22.18 1.344
w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$ 30.00 6.131 22.45 1.200 30.13 6.220 22.17 1.333
w/o $\mathrm{CA}_{\mathrm{diff}}^{\prime}$ 31.55 6.345 22.64 1.315 32.00 6.530 22.81 1.433
w/o Propagation 33.23 6.399 22.88 1.501 32.80 6.780 23.01 1.523
Ours 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
Figure 4: Visualization of mask strategy ablations. Quantitative results on Pick-a-Pic and DrawBench across SD2.1, SDXL, and FLUX for HPS, AES, Pick, and IR metrics. Compared variants include w/o Mask, Random Mask, w/o $\mathrm{CA}_{\mathrm{orig}}^{\prime}$, w/o $\mathrm{CA}_{\mathrm{diff}}^{\prime}$, and w/o Propagation. LoTTS consistently outperforms these ablated versions, highlighting the importance of each mask generation component.
D.2 Prompt Design
As in Tab. 4, this ablation investigates how different prompt formulations influence the effectiveness of LoTTS. We compare two strategies: (1) “a good/bad image of {p}”, and (2) “a high/low-quality image of {p}”, where {p} denotes the original text prompt. These two designs differ subtly in semantics: the former focuses on the overall desirability of the image, while the latter emphasizes rendering fidelity and fine-grained visual quality. Across all three model families (SD2.1, SDXL, and FLUX), both prompt strategies consistently improve generation quality, with only mild differences between them. For SD2.1 and FLUX, the two prompts perform similarly across all metrics, with the “high/low-quality” formulation yielding slightly higher Pick and IR scores. For SDXL, however, the “good/bad” prompt generally achieves better HPS and AES on Pick-a-Pic, suggesting that SDXL is more responsive to value-oriented descriptors than to explicit quality modifiers. Overall, the results indicate that LoTTS is robust to the specific choice of prompt template: both designs effectively provide the contrastive supervision needed for preference-aware optimization. The observed differences are small and model-specific, implying mild lexical sensitivity but no strong dependence on prompt wording.
Table 4: Ablation on Prompt Strategies for LoTTS. We evaluate two prompt designs, where i=0i{=}0 corresponds to “a good/bad image of {p}”, and i=1i{=}1 corresponds to “a high/low-quality image of {p}”, where {p} denotes the original text prompt. Results are reported on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model i{i} Pick-a-Pic DrawBench
[…]
SD2.1 0 24.23 5.821 21.31 0.679 23.27 5.923 21.33 0.701
1 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
SDXL 0 28.45 6.312 22.27 1.105 28.89 6.318 22.21 1.131
1 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
FLUX 0 33.11 6.505 23.00 1.604 33.78 6.889 23.15 1.620
1 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
Appendix E Parameter Analysis
E.1 Hyperparameter Sensitivity
We analyze the sensitivity of LoTTS to its key hyperparameters: number of refinements $k$, mask area ratio $r$, and noise injection step $t_{0}$. Tab. 5, Tab. 6, Tab. 7 and Fig. 5 summarize the results.
Number of refinements $k$.   Across SD2.1, SDXL, and FLUX, the best performance consistently occurs at $k=2$. In all experiments, we fix the final number of images per prompt at $N=9$, so changing $k$ redistributes this fixed sampling budget between the number of distinct initial samples and the number of local refinements applied to each of them. When $k$ increases, more budget is spent on local resampling around each starting point, but fewer distinct starting points are available. Although each local refinement is stochastic, the local generation space around a fixed starting point is inherently limited: additional local samples do not provide substantially better refinement candidates. In contrast, reducing the number of initial samples significantly weakens global diversity, which is often very important for discovering high-quality candidates under a fixed budget. With $k=2$, this balance is optimal: there are enough distinct initial samples to maintain global diversity, and enough local variations to explore meaningful refinement modes. In comparison, $k=1$ offers insufficient local exploration, while $k=8$ collapses global diversity by relying on a single starting point, leading to notably worse Best-of-9 performance. Thus, $k=2$ provides the most effective trade-off under a fixed sampling budget.
Table 5: Quantitative results w.r.t. number of refinements $k$. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model $k$ Pick-a-Pic DrawBench
[…]
SD2.1 1 23.50 5.781 21.20 0.581 22.23 5.799 21.24 0.540
2 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
8 21.34 5.367 20.43 0.363 21.90 5.678 20.38 0.354
SDXL 1 26.31 6.205 22.15 0.965 27.01 6.243 22.21 0.944
2 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
8 24.45 6.111 21.53 0.781 24.84 6.144 21.56 0.877
FLUX 1 31.46 6.445 22.89 1.550 33.28 6.631 22.95 1.457
2 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
8 30.54 6.308 22.57 1.244 30.81 6.332 22.15 1.200
Mask area ratio $r$.   Tab. 6 shows that a moderate mask area ratio (about $50\%$) consistently yields the best performance across SD2.1, SDXL, and FLUX. This trend reflects a fundamental trade-off between defect coverage and structural preservation. When the ratio is small (e.g., $r=20\%$), the mask typically captures only the most salient defect regions while ignoring secondary or subtle artifacts. As a result, many imperfections remain unrefined, lowering both the improvement rate (IR) and aesthetic gains (HPS, AES). This under-refinement behavior is evident in SD2.1, SDXL, and FLUX, where $r=20\%$ systematically underperforms the $50\%$ setting across nearly all metrics. Increasing the ratio to a moderate level ($r=50\%$) substantially improves all metrics. At this setting, the mask covers not only primary defects but also neighboring regions that contribute to local inconsistency, enabling more comprehensive correction. This leads to strong gains in HPS, AES, and IR across all three models. However, when the mask becomes too large ($r=80\%$), performance begins to decline. Excessively large masks start to overwrite high- quality regions, particularly detailed textures or well-formed structures, making the refinement process resemble global resampling. Interestingly, FLUX obtains the highest AES at $r=80\%$, but its HPS, Pick, and IR metrics do not surpass the $50\%$ setting, indicating that although aesthetic smoothness may improve, semantic fidelity and user preference do not benefit from excessive mask coverage. Overall, the optimal range around $50\%$ aligns with the observation that localized refinement should balance two competing goals: (1) covering enough area to correct both major and subtle defects, while (2) preserving globally coherent structure and avoiding unnecessary alteration of high-quality content. This trade-off consistently favors a moderate ratio across all tested models and datasets.
Table 6: Quantitative results w.r.t. mask area ratio $r$. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model $r$ Pick-a-Pic DrawBench
[…]
SD2.1 20 23.29 5.780 21.22 0.630 22.59 5.778 21.33 0.532
50 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
80 23.34 5.978 21.23 0.640 22.39 5.791 21.43 0.620
SDXL 20 26.13 6.244 22.15 0.895 27.46 6.243 22.13 0.894
50 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
80 27.03 6.256 22.13 0.985 26.98 6.313 22.21 0.944
FLUX 20 31.46 6.432 22.99 1.550 33.48 6.623 22.89 1.545
50 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
80 32.66 6.632 23.00 1.459 33.18 6.724 22.93 1.615
Noise injection step $t_{0}$.   Tab. 7 shows that the optimal noise injection step varies noticeably across SD2.1, SDXL, and FLUX, reflecting the different generative dynamics of their underlying architectures. For the U-Net–based diffusion models SD2.1 and SDXL, injecting noise at an early-to-mid denoising step produces the strongest results. Specifically, SD2.1 achieves its best performance at $t_{0}=25$ out of 50 steps, while SDXL peaks at $t_{0}=15$ out of 30 steps. At these intermediate stages, the model retains enough semantic structure from the original sample to preserve the global layout while still having sufficient freedom to update defect regions. This balance leads to clear gains in HPS, AES, and IR. The curves in Tab. 7 further show that both too-early injection ($t_{0}=0$) and too-late injection ($t_{0}=40$ for SD2.1, $t_{0}=25$ for SDXL) result in inferior performance: injecting noise too early disrupts global structure, while injecting too late limits the model’s ability to make meaningful local corrections. In contrast, FLUX exhibits a markedly different pattern due to its rectified flow architecture. Unlike U-Net diffusion, FLUX rapidly forms stable global structures at the beginning of its generation trajectory. As a result, even an extremely small noise step such as $t_{0}=1$ provides enough flexibility to refine local defects without disturbing the overall composition. This is reflected in its consistently strong performance at $t_{0}=1$, achieving the best HPS, AES, Pick, and IR across both datasets. Larger injection steps, however, degrade performance: injecting noise at $t_{0}=5$ or $8$ overwrites too much of the already-converged global structure, causing the refined image to deviate from the intended semantics. Taken together, these observations highlight that the optimal noise injection stage is tightly coupled to the generative process of each architecture. U-Net–based diffusion models favor early-to-mid stages where both structure and flexibility coexist, whereas rectified flow models like FLUX require only a minimal noise injection due to their fast convergence properties. LoTTS naturally adapts to these dynamics by allowing $t_{0}$ to control the balance between preserving global coherence and enabling local refinement.
Table 7: Quantitative results w.r.t. noise injection step $t_{0}$. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model $t_{0}$ Pick-a-Pic DrawBench
[…]
SD2.1 0 20.54 5.487 20.53 0.238 21.68 5.556 20.33 0.254
10 23.32 5.755 21.22 0.678 23.19 5.910 21.07 0.658
25 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
40 21.42 5.775 21.01 0.650 23.21 5.881 21.27 0.648
SDXL 0 23.81 6.041 21.34 0.683 23.95 6.054 21.14 0.667
5 25.13 6.254 22.17 1.062 28.75 6.301 21.08 1.051
15 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
25 27.30 6.194 22.24 1.002 28.65 6.283 21.88 1.091
FLUX 0 29.55 6.298 22.57 1.111 29.59 6.233 22.15 1.117
1 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
5 30.15 6.312 22.77 1.211 29.93 6.349 22.48 1.263
8 30.01 6.301 22.60 1.200 29.88 6.301 22.30 1.145
Visualization.   Fig. 5 visualizes the behavior of LoTTS under different hyperparameter settings. Across all models and metrics, the curves vary smoothly and exhibit clear optima, indicating that LoTTS is stable and well-conditioned. The best-performing points form broad, well-defined basins rather than isolated spikes, showing that LoTTS does not rely on fragile hyperparameter choices and is robust across a wide operating range.
Figure 5: Hyperparameter analysis of LoTTS. Performance on Pick-a-Pic and DrawBench with respect to (top) the number of refinements $k$, (middle) mask area ratio $r$, and (bottom) noise injection step $t_{0}$, across SD2.1, SDXL, and FLUX. Stars mark the best-performing configurations.
E.2 Scaling Comparison with Best-of-$N$
We evaluate how LoTTS scales with the total number of generated images $N$ and compare it against a pure global Best-of-$N$ baseline. In both cases, $N$ represents the total number of images generated per prompt. For Best-of-$N$, this is simply the number of globally sampled candidates. For LoTTS, however, the same budget $N$ is redistributed between the number of initial samples and the number of local refinements performed around each initial sample. For example, $N=1$ corresponds to pure global sampling, $N=3$ corresponds to $1$ initial sample with $k+1=3$ local candidates, $N=6$ corresponds to $2$ initial samples with $k+1=3$ refinements each, and $N=9$ corresponds to the default LoTTS setting of $3$ initial samples with $k+1=3$ refinements each.
Scaling behavior of LoTTS.   Tab. 8 shows that LoTTS exhibits smooth and consistent performance gains as the total number of generated images $N$ increases from 1 to 9. Even with extremely small budgets, such as $N=3$, LoTTS already achieves notable improvements over pure global sampling ($N=1$). For example, on SD2.1, HPS increases from 20.44 at $N=1$ to 21.12 at $N=3$, accompanied by an IR gain from 0.236 to 0.413. As $N$ continues increasing, these gains become more substantial: at $N=6$, SD2.1 reaches HPS 22.29 and IR 0.531; and at $N=9$, LoTTS achieves its strongest performance with HPS 24.52 and IR 0.680. A similar pattern appears in SDXL and FLUX. On SDXL, HPS improves from 23.44 ($N=1$) → 25.23 ($N=3$) → 26.21 ($N=6$) → 28.23 ($N=9$), with IR rising from 0.680 to 1.102. FLUX shows even stronger scaling: HPS increases from 29.34 ($N=1$) to 31.23 ($N=3$), then to 32.24 ($N=6$), and ultimately reaches 33.33 at $N=9$. The IR metric follows the same trend, improving from 1.038 to 1.605. These consistent improvements across all models and metrics indicate that each additional global starting point introduced by a larger budget $N$ provides new semantic modes for LoTTS to refine. Since every starting point is further expanded into multiple locally refined variants, LoTTS is able to generate a rich set of high-quality candidates around these diverse global modes.
Table 8: Quantitative results of LoTTS w.r.t. sample count NN. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model NN Pick-a-Pic DrawBench
[…]
SD2.1 1 20.44 5.377 20.32 0.236 21.34 5.546 20.23 0.244
3 21.12 5.532 20.56 0.413 21.88 5.623 20.54 0.398
6 22.29 5.712 20.99 0.531 21.45 5.807 21.00 0.591
9 24.52 5.805 21.32 0.680 23.29 5.911 21.47 0.698
SDXL 1 23.44 6.011 21.18 0.680 23.84 6.034 21.09 0.657
3 25.23 6.124 21.55 0.712 25.11 6.123 21.55 0.701
6 26.21 6.153 21.89 0.813 26.23 6.241 21.98 0.812
9 28.23 6.304 22.30 1.102 28.90 6.321 22.38 1.111
FLUX 1 29.34 6.298 22.07 1.038 29.28 6.223 22.05 1.100
3 31.23 6.370 22.53 1.203 30.45 6.350 22.47 1.321
6 32.24 6.434 22.88 1.523 32.23 6.591 23.01 1.521
9 33.33 6.501 23.04 1.605 33.90 6.890 23.21 1.623
Scaling behavior of Best-of-$N$.   Tab. 9 shows that the Best-of-$N$ baseline consistently improves as the number of globally sampled images increases. This trend is expected: drawing more samples enlarges the global search space, increasing the likelihood of encountering rare but high-quality generations. As $N$ grows, the tails of the model’s output distribution are explored more thoroughly, naturally raising the probability of selecting an image with stronger aesthetics or semantic alignment. The numerical results reflect this behavior clearly. For SD2.1, HPS improves from 20.44 at N=1N{=}1 to 21.56 at N=9N{=}9, then to 22.21 at N=17N{=}17, ultimately reaching 24.23 at N=25N{=}25. SDXL follows a similar progression, increasing from 23.44 to 25.44, then to 26.02, and finally to 28.21 when N=30N{=}30. FLUX exhibits the same scaling pattern, with HPS rising from 29.34 at N=1N{=}1 to 30.23 at N=9N{=}9, then to 31.14 at N=14N{=}14, and ultimately reaching 33.32 at N=36N{=}36. The other metrics—AES, Pick, and IR—show comparable monotonic gains, though with diminishing increments as $N$ becomes large. Overall, these results confirm that Best-of-$N$ behaves like a global search procedure: as more samples are drawn, the chance of finding a superior candidate increases, leading to smooth and predictable scaling trends across models and evaluation metrics.
Table 9: Quantitative results of Best-of-NN baseline w.r.t. sample count NN. Results on Pick-a-Pic and DrawBench for SD2.1, SDXL, and FLUX.
Model NN Pick-a-Pic DrawBench
[…]
SD2.1 1 20.44 5.377 20.32 0.236 21.34 5.546 20.23 0.244
9 21.56 5.534 21.04 0.470 22.45 5.589 20.59 0.446
17 22.21 5.755 21.11 0.524 25.17 5.817 21.10 0.602
25 24.23 5.825 21.26 0.681 23.33 5.915 21.50 0.701
SDXL 1 23.44 6.011 21.18 0.680 23.84 6.034 21.09 0.657
9 25.44 6.198 22.01 0.790 25.27 6.238 22.03 0.756
20 26.02 6.201 22.11 0.833 26.31 6.233 22.30 0.832
30 28.21 6.302 22.21 1.108 28.88 6.330 22.42 1.113
FLUX 1 29.34 6.298 22.07 1.038 29.28 6.223 22.05 1.100
9 30.23 6.299 22.89 1.235 30.46 6.290 22.23 1.221
14 31.14 6.343 22.93 1.521 32.90 6.599 22.98 1.533
36 33.32 6.531 23.09 1.615 33.88 6.891 23.22 1.620
[…]
Appendix F Discussion
This supplementary discussion expands upon the brief analysis in Sec. 5 and provides a unified perspective connecting empirical failure modes (illustrated in Fig. 6) with the theoretical guarantees and boundary conditions formalized in Appendix G. We organize the discussion into two parts: Failure Analysis and Directions for Improvement.
F.1 Failure Analysis
Imperfect masks and unstable attention signals.   A major source of failure arises from attention-based masks missing true defects or selecting unstable high-frequency regions. Subtle artifacts, micro-textures, or weak structural inconsistencies may receive insufficient attention, leading to low recall, while noisy activations near boundaries often generate false positives. These behaviors correspond directly to the theoretical sensitivity to precision and recall discussed in Appendix G. In particular, insufficient recall violates the requirement in  equation G.3.1, preventing LoTTS from reaching enough defective patches to yield meaningful improvement. Conversely, poor precision inflates the false-positive harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$, which must remain smaller than $q\overline{\delta}$ for each selected patch to provide positive expected benefit, as formalized in  equation 22.
Global defects and violation of sparsity.   A second category of failures arises from global semantic or geometric errors, such as incorrect perspective, inconsistent lighting, or cross-object misalignment. Because LoTTS modifies only a limited set of selected patches, it lacks the ability to enact coordinated, image-wide corrections, and therefore cannot address defects that require globally coherent changes. Such cases fundamentally violate the sparse-defect assumption $s\ll M$ that underpins Theorem 1 and Corollary 3. When defects behave as if they are dense or spatially widespread, the global-harm amplification factor $(\tfrac{M}{s}-1)h_{g}\overline{\gamma}$ diminishes, thereby eliminating the structural advantage that makes localized refinement more compute-efficient than global resampling. This explains why global TTS can occasionally succeed in scenes with strong global inconsistencies, despite its disadvantages in sparse-defect settings.
Inputs at quality extremes.   LoTTS also encounters difficulties when the input image lies at the extremes of the quality spectrum. In very low-quality images, the effective set of defects behaves as though $s\approx M$, which violates the sparsity regime and limits the potential for safe and effective localized improvement. Conversely, when the input image is already near-perfect, the expected repair gain $\overline{\delta}$ becomes extremely small, reducing the net improvement margin and making the refinement process susceptible to small perturbation artifacts. Both scenarios follow directly from Theorem 1, which predicts that LoTTS loses its advantage either when the intrinsic repair benefit collapses or when the sparsity assumption no longer holds.
Failure summary.   Across these categories, the observed empirical failures align closely with the theoretical failure conditions formalized in Corollary 5. When defects are dense or global, the global-harm asymmetry disappears and LoTTS no longer enjoys a structural advantage. When mask precision is extremely low, false-positive edits impose excessive local harm that overwhelms the expected benefit. And when the local repair mechanism yields only a limited improvement relative to $\theta_{g}$, localized refinement lacks sufficient corrective strength. These patterns demonstrate that LoTTS’s limitations arise not from unpredictable behavior, but from predictable breakdowns of the assumptions required for theoretical dominance.
Figure 6: Failure cases of LoTTS. While LoTTS reduces many local artifacts, a few failures remain in boundary geometry, in scenes requiring global coherence or containing low-texture regions (e.g., sky, court floor), and in cases with only subtle defects where refinements bring limited improvements.
F.2 Future Directions
Restrict computation strictly within the mask.   A promising direction is to integrate algorithms whose computation is explicitly confined to the mask region, such as mask-restricted resampling. By design, these methods allocate compute exclusively to the regions that require refinement, effectively realizing an extreme form of cost asymmetry ($C_{\ell}\ll C_{g}$). According to Corollary 2, this places the method directly in a regime where localized refinement is provably more compute-efficient than global resampling. Such mask-restricted algorithms therefore offer a principled pathway toward significantly improving test-time scaling efficiency: the same compute budget yields far more effective updates simply because no computation is wasted on clean regions.
Baseline beyond Best-of-$N$.   Although our current implementation adopts a Best-of-$N$ sampling scheme within the masked region, this strategy represents only the simplest form of local search. A more powerful alternative is to integrate global search algorithms—such as beam search—that branch along the denoising trajectory, but do so only within the masked area. Such mask-restricted branching preserves the locality and compute efficiency of LoTTS while enabling a substantially richer exploration of refinement hypotheses than Best-of-$N$. This opens the door to “global search applied locally,” where complex trajectory-level decisions are made without ever incurring the cost or harm of full-image resampling.
Early-stop mechanisms for low-potential cases.   Another practical extension is to incorporate an early-stop criterion that detects when an image has little potential for localized improvement. In many cases, the mask may identify only marginal or low-confidence regions, or the estimated patch-level quality signals may indicate that no meaningful refinement can be achieved within the masked area. Continuing localized resampling in such situations not only yields negligible benefit but also risks introducing unnecessary perturbations. An early-stop mechanism—triggered by low mask coverage, low defect confidence, or minimal improvement predicted from intermediate diffusion states—would allow the system to gracefully abstain from refinement when appropriate. This ensures that compute is spent only when LoTTS is theoretically capable of providing positive expected gain, further improving test-time efficiency and robustness.
Improving mask-generation strategies.   The current mask is constructed from relative internal inconsistencies, which can miss globally subtle or low-contrast defects. A more principled alternative is to incorporate absolute quality cues, such as no-reference IQA metrics, to better identify patches with genuine improvement potential. In addition, richer mask-generation strategies based on multi-signal fusion (e.g., trajectory disagreement, latent-gradient saliency) may significantly improve both precision and recall. By stabilizing the mask and grounding it in multiple complementary signals, these approaches can move the mask into the regime where LoTTS is theoretically guaranteed to yield positive expected gain ( equation 22), thereby enhancing robustness and overall refinement effectiveness.
Appendix G Theoretical Analysis
G.1 Problem Setup and Preliminaries
Overview.   Our theoretical analysis proceeds in three steps. (1) We first express global image quality through a patch-level lower bound, which allows local refinements to be analyzed independently. (2) We then characterize the stochastic behavior of the mask via precision and recall, which determines how often LoTTS touches defective or clean patches. (3) Combining these two components, we compute the expected gain of global and local TTS and derive the condition under which localized refinement is more compute-efficient.
Notation.   We analyze the conditions under which localized TTS achieves higher quality or efficiency than global TTS. For clarity, Tab. 10 summarizes the variables used throughout this section. This ensures every symbol in the derivation has an explicit meaning.
Symbol Description
$M$ total number of non-overlapping image patches
$s$ number of defective patches ($s\ll M$)
$S$ set of defective patches, with $|S|=s$
$\widehat{S}$ set of patches selected by LoTTS for resampling
$\pi$ precision: fraction of selected patches that are truly defective
$\rho$ recall: fraction of defective patches successfully selected
$x$ generated image; $x_{j}$ denotes patch $j$
$r(x)$ perceptual quality functional of image $x$
$r_{j}(x_{j})$ patch-level quality score
$w_{j}$ importance weight of patch $j$ in quality lower bound
$\delta_{j}$ expected gain if defective patch $j$ is repaired
$\gamma_{j}$ expected loss if clean patch $j$ is harmed
$\overline{\delta}$ average weighted repair gain over defective patches
$\overline{\gamma}$ average weighted harm over clean patches
$\theta_{g}$ probability of repairing a defective patch under global resampling
$q$ probability of repairing a defective patch under localized resampling
$h_{g}$ probability of harming a clean patch under global resampling
$h_{\ell}$ probability of harming a clean patch under localized resampling
$C_{g}$ compute cost of one global resampling trial
$C_{\ell}$ compute cost of one localized resampling trial
$B$ total compute budget (e.g., measured in NFEs)
Table 10: Notation used in the theoretical analysis.
G.1.1 Image Quality Decomposition
Before comparing global and local strategies, we first relate global quality $r(x)$ to patch-level scores. To lower bound global improvement by local improvements, we adopt a standard additive inequality:
[…]
where $w_{j}\geq 0$ are weights reflecting importance of each patch. This allows us to measure global gain through weighted patch-level gains.
G.1.2 Mask Selection Statistics
We next connect the mask’s precision/recall to expected true/false selections. This step is necessary to compute how many patches are truly improved versus unnecessarily modified.
Lemma 1 (Expected TP/FP Statistics).
[…]
Proof.
By the definition of recall $\rho=\frac{\mathbb{E}[\mathrm{TP}]}{s}$, we immediately have:
$\mathbb{E}[\mathrm{TP}]=\rho\,s.$ (12)
By the (operational) definition of precision $\pi=\frac{\mathbb{E}[\mathrm{TP}]}{\mathbb{E}[|\widehat{S}|]}$, we get:
$\mathbb{E}[|\widehat{S}|]=\frac{\mathbb{E}[\mathrm{TP}]}{\pi}=\frac{\rho\,s}{\pi}.$ (13)
Finally, using linearity of expectation and $\mathrm{FP}=|\widehat{S}|-\mathrm{TP}$,
$\mathbb{E}[\mathrm{FP}]=\mathbb{E}[|\widehat{S}|]-\mathbb{E}[\mathrm{TP}]=\frac{\rho\,s}{\pi}-\rho\,s=\rho\,s\Big(\tfrac{1}{\pi}-1\Big).$ (14)
Discussion.
  • Why use $\pi=\frac{\mathbb{E}[\mathrm{TP}]}{\mathbb{E}[|\widehat{S}|]}$? Precision is often defined as $\mathbb{E}\!\left[\frac{\mathrm{TP}}{|\widehat{S}|}\mathbf{1}\{|\widehat{S}|>0\}\right]$. To avoid division-by-zero and keep algebra tractable, we adopt the ratio-of-expectations form, which is standard in compute-budget analyses. If one insists on the expectation-of-ratio definition, Jensen-type arguments yield the bound $\mathbb{E}[|\widehat{S}|]\geq\frac{\mathbb{E}[\mathrm{TP}]}{\pi}$, so our equalities become tight upper/lower bounds; the conclusions below only change by harmless inequalities.
[…]
  • What randomness is averaged over? Expectations are taken over the randomness of mask construction (and, if applicable, sampling noise). In practice, the mask may be stochastic due to thresholding on noisy attention maps, variation across diffusion trajectories, or random perturbations used during test-time evaluation. No independence assumptions are required; linearity of expectation suffices.
G.2 Main Theorem
This section derives the expected improvement of a single global or localized resampling step, and then extends it to a fixed compute budget. These expressions form the basis of the dominance condition analyzed in the next subsection.
G.2.1 Expected Gain per Resampling Step
We begin by analyzing the expected gain of a single resampling trial. Let $x$ denote the image before the update and $x^{+}$ the image after the update, and define
$\Delta r:=r(x^{+})-r(x).$ (15)
To relate global quality to patch-level behavior, we use the standard additive lower bound:
[…]
where $w_{j}\geq 0$ reflects the importance of patch $j$. This inequality allows us to lower-bound the global improvement through weighted patch-level changes:
$\Delta r\ \geq\ \sum_{j=1}^{M}w_{j}\big(r_{j}(x_{j}^{+})-r_{j}(x_{j})\big).$ (17)
For each defective patch $j\in S$, let $\delta_{j}\geq 0$ denote the expected improvement when the patch is successfully repaired; for each clean patch $j\notin S$, let $\gamma_{j}\geq 0$ denote the expected loss when the patch is harmed. Define the weighted averages
$\overline{\delta}:=\frac{1}{s}\sum_{j\in S}w_{j}\delta_{j},\hskip 28.80008pt\overline{\gamma}:=\frac{1}{M-s}\sum_{j\notin S}w_{j}\gamma_{j},$ (18)
where $s=|S|\ll M$ is the number of defective patches. These quantities do not require independence across patches; they capture only the expected per-patch effect of one resampling step.
Global resampling.   In global test-time scaling, all patches are resampled. Each defective patch $j\in S$ is repaired with probability $\theta_{g}$, and each clean patch $j\notin S$ is harmed with probability $h_{g}$. Aggregating over all patches gives the per-trial expected improvement:
$\mathbb{E}[\Delta r]_{\mathrm{global}}\ \geq\ s\,\theta_{g}\,\overline{\delta}\;-\;(M-s)\,h_{g}\,\overline{\gamma}.$ (19)
The benefit scales with the $s$ defective patches, while the harm scales with the much larger number (Ms)(M-s) of clean patches.
Localized resampling (LoTTS).   In localized test-time scaling, only patches selected by the mask $\widehat{S}$ are resampled: potentially repairing true positives $S\cap\widehat{S}$ and harming false positives $\widehat{S}\setminus S$.
Let recall $\rho$ and precision $\pi$ be defined as in Lemma 1. Let $q$ be the probability of repairing a defective selected patch, and $h_{\ell}$ the probability of harming a selected clean patch. Using Lemma 1, the expected numbers of selected defective and clean patches are $\rho s$ and $\rho s(\tfrac{1}{\pi}-1)$, respectively, leading to
$\mathbb{E}[\Delta r]_{\mathrm{local}}\ \geq\ \rho\,s\,q\,\overline{\delta}\;-\;\rho\,s\Big(\tfrac{1}{\pi}-1\Big)\,h_{\ell}\,\overline{\gamma}.$ (20)
Interpretation.   Global resampling can repair all $s$ defective patches, but pays a harm penalty scaling with the large number (Ms)(M-s) of clean patches. LoTTS repairs only $\rho s$ defective patches in expectation, but its harm penalty grows only with the false positives, of order $\rho s(\tfrac{1}{\pi}-1)$ rather than $M$. This asymmetry is the essence of the sparse-defect advantage: when $s\ll M$, localized refinement focuses compute on problematic regions while avoiding widespread degradation.
G.2.2 Expected Gain per Compute Unit
To compare the two strategies under a fixed compute budget $B$, we normalize their per-trial gains by their respective costs $C_{g}$ and $C_{\ell}$. Since a total of $B/C_{g}$ global trials or $B/C_{\ell}$ local trials can be performed, we obtain:
Theorem 1 (Expected Quality Gains).
For a total compute budget $B$, let $C_{g}$ and $C_{\ell}$ denote the cost of a single global and localized trial, and let $\overline{\delta}$ and $\overline{\gamma}$ denote the average per–patch repair gain and harm penalty. Then the expected quality improvements of Global TTS and LoTTS satisfy: $\boxed{\begin{aligned} \mathbb{E}[\Delta r]_{\mathrm{global},B}&\;\geq\;\frac{B}{C_{g}}\Big[\,s\,\theta_{g}\,\overline{\delta}\;-\;(M-s)\,h_{g}\,\overline{\gamma}\,\Big],\\[3.0pt] \mathbb{E}[\Delta r]_{\mathrm{local},B}&\;\geq\;\frac{B}{C_{\ell}}\Big[\,\rho s\,q\,\overline{\delta}\;-\;\rho s\!\big(\tfrac{1}{\pi}-1\big)h_{\ell}\overline{\gamma}\,\Big].\end{aligned}}$
Interpretation.   Both methods yield an expected quality change equal to (average repair gain) minus (average harm penalty), scaled by the number of trials affordable under budget $B$.
These compute-normalized expressions provide the basis for the dominance condition analyzed in the next subsection.
G.3 Derived Corollaries
Having established the expected-gain expressions, We now determine when localized resampling (LoTTS) achieves a higher compute-normalized expected improvement than global test-time scaling.
G.3.1 General Corollaries
Corollary 1 (General Case).
LoTTS achieves higher compute-normalized expected improvement than Global TTS if $\boxed{\begin{gathered}\frac{\rho}{C_{\ell}}\Big(q\overline{\delta}-(\tfrac{1}{\pi}-1)\,h_{\ell}\overline{\gamma}\Big)\\[6.0pt] >\;\frac{1}{C_{g}}\Big(\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)\,h_{g}\overline{\gamma}\Big)\end{gathered}}$
Interpretation.   The left-hand side represents the compute-normalized expected improvement contributed by the patches selected by LoTTS: recall $\rho$ determines how many defective patches are reached, $q\overline{\delta}$ is the average benefit of repairing a defect, and $(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ accounts for harm from false positives. The right-hand side describes the corresponding global gain, where the potential harm is amplified by (Ms)(M-s) clean patches rather than by the (much smaller) number of false positives.
The dominance inequality tells us when LoTTS is better, but it does not yet answer the key practical question: how accurate must the mask be for LoTTS to actually win? This subsection makes this explicit by deriving the required recall $\rho$ and precision $\pi$ thresholds.
Required recall.   Solving the inequality in Corollary 1 for $\rho$ yields
$\rho\ >\ \frac{C_{\ell}}{C_{g}}\,\frac{\theta_{g}\overline{\delta}-\big(\tfrac{M}{s}-1\big)h_{g}\overline{\gamma}}{q\overline{\delta}-(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}}$
$\text{provided}\quad q\overline{\delta}-(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}>0.$ (21)
The denominator is the expected net benefit of selecting a single patch; it must be positive for localized refinement to be useful.
Required precision.   A convenient sufficient condition for the denominator of equation G.3.1 to be positive is
$q\overline{\delta}-(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;>\;0\qquad\Longleftrightarrow\qquad\pi\;>\;\frac{1}{\,1+\frac{q\overline{\delta}}{h_{\ell}\overline{\gamma}}\,}.$ (22)
This threshold is easier to satisfy when true repairs provide large gains ($q\overline{\delta}$ high) or when false-positive edits cause limited harm ($h_{\ell}\overline{\gamma}$ low). In either case, even imperfect masks yield positive expected gain per selected patch.
G.3.2 Specialized Corollaries
We now present simplified dominance conditions obtained from Theorem 1 under practically relevant regimes. Each corollary corresponds directly to one of the scenarios in the Q&A of Section G.4.
Corollary 2 (Dominance Under Cost Asymmetry).
From Theorem 1, LoTTS dominates global TTS iff $\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;>\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$ (23) In the regime where $C_{\ell}\ll C_{g},\hskip 28.80008pt\Delta_{\ell}^{\left(1\right)}=\Theta(\Delta_{g}^{\left(1\right)})>0,$ condition equation 23 simplifies to $\boxed{\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;\gg\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.}$ Thus LoTTS achieves strictly larger compute-normalized gain under any fixed compute budget $B$.
Proof.
Step 1: Substitute the expressions from Theorem 1.   The expected improvements under total budget $B$ are
$\mathbb{E}[\Delta r]_{\mathrm{local},B}=\frac{B}{C_{\ell}}\,\Delta_{\ell}^{\left(1\right)},\hskip 28.80008pt\mathbb{E}[\Delta r]_{\mathrm{global},B}=\frac{B}{C_{g}}\,\Delta_{g}^{\left(1\right)}.$
Step 2: Translate dominance into an inequality.   LoTTS dominates global TTS exactly when
$\mathbb{E}[\Delta r]_{\mathrm{local},B}\;>\;\mathbb{E}[\Delta r]_{\mathrm{global},B},$
which becomes
$\frac{B}{C_{\ell}}\,\Delta_{\ell}^{\left(1\right)}\;>\;\frac{B}{C_{g}}\,\Delta_{g}^{\left(1\right)}.$
Step 3: Cancel the common positive factor $B$.   Since $B>0$, inequality equation G.3.2 is equivalent to
$\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;>\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}},$
which is the exact dominance condition in the corollary.
Step 4: Apply the cost-asymmetry assumptions.   Assume
$C_{\ell}\ll C_{g},\hskip 28.80008pt\Delta_{\ell}^{\left(1\right)}\approx\Delta_{g}^{\left(1\right)}>0.$
The first assumption implies
$\frac{1}{C_{\ell}}\gg\frac{1}{C_{g}},$
and the second implies that $\Delta_{\ell}^{\left(1\right)}$ and $\Delta_{g}^{\left(1\right)}$ are of comparable magnitude.
Step 5: Compare the compute-normalized benefits.   Using these relations,
$\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}=\Delta_{\ell}^{\left(1\right)}\frac{1}{C_{\ell}}\approx\Delta_{g}^{\left(1\right)}\frac{1}{C_{\ell}}\;\gg\;\Delta_{g}^{\left(1\right)}\frac{1}{C_{g}}=\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$
Thus
$\frac{\Delta_{\ell}^{\left(1\right)}}{C_{\ell}}\;\gg\;\frac{\Delta_{g}^{\left(1\right)}}{C_{g}}.$
Step 6: Conclude dominance for any fixed budget $B$.   Substituting this back into equation G.3.2 shows that
$\mathbb{E}[\Delta r]_{\mathrm{local},B}>\mathbb{E}[\Delta r]_{\mathrm{global},B},$
for any fixed $B$, establishing the corollary. ∎
Interpretation.   When local sampling is substantially cheaper, compute becomes the decisive advantage: LoTTS performs many more trials within the same budget, so even moderate $(\rho,\pi,q)$ are sufficient for dominance.
Corollary 3 (Dominance Under Equal-Cost Sparse Regime).
Assume $C_{\ell}=C_{g}$. Then Theorem 1 gives the exact condition $(\rho q-\theta_{g})\overline{\delta}\;>\;\Big(\tfrac{M}{s}-1\Big)h_{g}\overline{\gamma}+\rho\Big(\tfrac{1}{\pi}-1\Big)h_{\ell}\overline{\gamma}.$ (23) In the sparse-defect regime $s\ll M$ and with benign local edits $h_{\ell}\ll h_{g}$, the dominant terms yield the approximation $\boxed{\rho q\;\gtrsim\;\theta_{g}+\frac{M}{s}\,\frac{h_{g}\overline{\gamma}}{\overline{\delta}}.}$ Thus, under equal compute, dominance is governed by the relative scaling of the harm terms.
Proof.
We begin from the exact dominance inequality in Theorem 1:
$\frac{B}{C_{\ell}}\Big(\rho q\overline{\delta}-\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\Big)\;>\;\frac{B}{C_{g}}\Big(\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big).$ (24)
Step 1: Specialize to $C_{\ell}=C_{g}$.   Under the equal-cost assumption $C_{\ell}=C_{g}$, the prefactor $B/C_{\ell}$ is strictly positive and common to both sides. Therefore we may cancel it from equation 24, yielding the exact dominance condition
$\rho q\overline{\delta}-\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;>\;\theta_{g}\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}.$ (25)
Step 2: Collect all repair terms on the left and all harm terms on the right.   Rearranging equation 25 gives
$(\rho q-\theta_{g})\overline{\delta}$ $>\;(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\;+\;\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}.$ (26)
equation 26 is algebraically equivalent to the original dominance condition and valid without approximation.
Step 3: Apply sparse-defect assumption.   When $s\ll M$, we have
$\tfrac{M}{s}-1=\Theta(\tfrac{M}{s}),$
so the global-harm term becomes
$(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\approx\frac{M}{s}h_{g}\overline{\gamma},$
which dominates all global contributions because it scales with the number of clean patches.
Step 4: Apply benign-local-edit assumption.   If $h_{\ell}\ll h_{g}$, then the local-harm term
$\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$
is negligible compared to $\frac{M}{s}h_{g}\overline{\gamma}$, since (i) it does not scale with $M$, and (ii) $h_{\ell}\ll h_{g}$ suppresses its magnitude.
Thus, in the joint regime $(s\ll M)$ and $(h_{\ell}\ll h_{g})$, the right-hand side of equation 26 satisfies the asymptotic approximation
$(\tfrac{M}{s}-1)h_{g}\overline{\gamma}+\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}\;\approx\;\frac{M}{s}h_{g}\overline{\gamma}.$
Step 5: Divide both sides by $\overline{\delta}$.   Since $\overline{\delta}>0$, we divide equation 26 to obtain
$\rho q-\theta_{g}\;\gtrsim\;\frac{M}{s}\frac{h_{g}\overline{\gamma}}{\overline{\delta}},$
or equivalently
$\rho q\;\gtrsim\;\theta_{g}+\frac{M}{s}\frac{h_{g}\overline{\gamma}}{\overline{\delta}}.$
This is exactly the simplified condition stated in the corollary. ∎
Interpretation.   When compute cost is equal, the repair gains of the two methods are comparable; the decisive asymmetry lies in the harm terms. Global harm scales with all (Ms)(M-s) clean patches, whereas local harm scales only with false positives. In sparse regimes this makes the global harm term much larger, allowing LoTTS to dominate even with moderate $(\rho,\pi,q)$.
Corollary 4 (Saturation of Best-of-NN Global Resampling).
For global Best-of-NN, $C_{g}(N)=N\,C_{g}\left(1\right),\hskip 28.80008pt\theta_{g}(N)=1-(1-\theta_{g}\left(1\right))^{N}.$ The compute-normalized improvement becomes $\boxed{\frac{1}{NC_{g}\left(1\right)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big).}$ Since $\theta_{g}(N)$ increases sublinearly in NN while the harm term remains unchanged, global improvement saturates and cannot overcome the harm term by increasing $N$.
Proof.
Step 1: Define the Best-of-$N$ procedure and its cost.   A Best-of-$N$ trial draws $N$ independent global resampling samples and selects the best one. Since one global draw costs $C_{g}\left(1\right)$, performing $N$ draws costs
$C_{g}(N)=N\,C_{g}\left(1\right).$
Step 2: Compute the defect repair probability.   Let $\theta_{g}\left(1\right)$ be the probability that a single global draw repairs a defect. Under independence across draws, the probability that none of the $N$ draws repairs the defect is $(1-\theta_{g}\left(1\right))^{N}$. Thus the probability that at least one draw repairs it is
$\theta_{g}(N)=1-\big(1-\theta_{g}\left(1\right)\big)^{N}.$
Step 3: Substitute into the global gain expression.   Plugging $C_{g}(N)$ and $\theta_{g}(N)$ into the global expression in Theorem 1, the expected improvement under budget $B$ becomes
$\mathbb{E}[\Delta r]_{\mathrm{global},B}=\frac{B}{C_{g}(N)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big).$
Step 4: Normalize by compute.   Dividing both sides by $B$ yields
$\frac{1}{C_{g}(N)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big)$
$=$ $\frac{1}{NC_{g}\left(1\right)}\Big(\theta_{g}(N)\overline{\delta}-(\tfrac{M}{s}-1)h_{g}\overline{\gamma}\Big),$
which is exactly the boxed expression in the corollary.
Step 5: Show saturation as $N$ increases.   The function
$\theta_{g}\left(N\right)=1-\left(1-\theta_{g}\left(1\right)\right)^{N}$
is increasing and concave in $N$. Indeed, its incremental increase satisfies
$\theta_{g}(N+1)-\theta_{g}(N)=(1-\theta_{g}\left(1\right))^{N}\,\theta_{g}\left(1\right),$
which is positive but strictly decreasing in $N$. Thus $\theta_{g}(N)$ approaches $1$ with diminishing returns.
Step 6: Conclude saturation of compute-normalized global gain.   The compute-normalized gain contains the prefactor $1/(NC_{g}\left(1\right))$, which decreases linearly in $N$, while the harm term
$(\tfrac{M}{s}-1)h_{g}\overline{\gamma}$
does not depend on $N$ at all. Therefore, as $N$ increases:
  • the gain term increases sublinearly (diminishing returns),
  • the harm term remains fixed,
  • the factor $1/(NC_{g}\left(1\right))$ shrinks linearly.
Consequently, the compute-normalized global improvement saturates and may eventually decline with increasing $N$. Unlike Best-of-$N$, LoTTS avoids global harm structurally, giving it an advantage that increasing $N$ cannot overcome.
Interpretation.   Increasing $N$ improves the chance of repairing a defect but does not reduce global harm. LoTTS avoids global harm structurally, giving it an advantage that Best-of-$N$ cannot eliminate.
Corollary 5 (Failure Conditions).
LoTTS fails to dominate whenever any one of the following conditions holds. These are exactly the scenarios where the dominance inequality in Theorem 1 reverses: Defects become dense ($s\!\approx\!M$), eliminating the sparse-defect advantage that normally suppresses global harm. Mask precision becomes extremely low, causing the false-positive harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ to dominate and outweigh the repair gains. Local repair probability is significantly smaller than the global one ($q\!\ll\!\theta_{g}$), reducing the expected improvement contributed by localized refinement.
Proof.
Step 1: Express dominance and its negation.   By Theorem 1, LoTTS dominates exactly when
$\mathbb{E}[\Delta r]_{\mathrm{local},B}>\mathbb{E}[\Delta r]_{\mathrm{global},B}.$
Thus LoTTS fails to dominate if and only if
$\mathbb{E}[\Delta r]_{\mathrm{local},B}\leq\mathbb{E}[\Delta r]_{\mathrm{global},B}.$
Step 2: Substitute the expressions from Theorem 1.   Using the explicit formulas for the expected improvements and cancelling the common positive factors yields
$\rho q\,\overline{\delta}-\rho\Big(\tfrac{1}{\pi}-1\Big)h_{\ell}\overline{\gamma}\;\leq\;\theta_{g}\,\overline{\delta}-\big(\tfrac{M}{s}-1\big)h_{g}\overline{\gamma}.$
Step 3: Group gain terms and harm terms.   Rearranging the inequality gives
$(\rho q-\theta_{g})\,\overline{\delta}\;\leq\;\Big(\tfrac{M}{s}-1\Big)h_{g}\overline{\gamma}+\rho\Big(\tfrac{1}{\pi}-1\Big)h_{\ell}\overline{\gamma}.$
Thus LoTTS fails precisely when the net local gain is no larger than the sum of the global harm and local false-positive harm.
Step 4: Interpret when the inequality can realistically hold.
  • Dense defects ($s\approx M$). When $s\approx M$, the factor $\tfrac{M}{s}-1$ becomes small. Thus the global harm term $(\tfrac{M}{s}-1)h_{g}\overline{\gamma}$ shrinks, eliminating the sparse-defect advantage. Global resampling no longer risks harming many clean patches, so the right-hand side can easily exceed the left-hand side.
  • Extremely low precision ($\pi\ll 1$). If precision is very low, $(\tfrac{1}{\pi}-1)$ becomes large. Then the local harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$ may dominate even when $h_{\ell}$ is small, causing LoTTS to apply many harmful edits to clean patches and making inequality equation G.3.2 hold.
  • Poor local repair probability ($q\ll\theta_{g}$). If $q$ is much smaller than $\theta_{g}$, then $\rho q-\theta_{g}$ is negative or strongly negative, making the left-hand side small even before considering harm. In such cases global resampling offers significantly greater repair gain, and LoTTS cannot compensate.
Step 5: Conclude the characterization of failure.   Each of the scenarios above corresponds to a way the right-hand side of equation G.3.2 can exceed the left-hand side: (i) global harm becomes small, (ii) local false-positive harm becomes large, (iii) local repair gain becomes too weak. These are precisely the regimes where
$\mathbb{E}[\Delta r]_{\mathrm{local},B}\leq\mathbb{E}[\Delta r]_{\mathrm{global},B},$
i.e. LoTTS fails to dominate.
Interpretation.   These corner cases correspond exactly to the settings in which LoTTS’s structural advantages—locality and sparse harm—no longer hold.
G.4 Practical Q&A
Theorem 1 and Corollary 1 specify when LoTTS is expected to outperform Global TTS. Below we summarize the main implications in a concise question–answer format, emphasizing why LoTTS is reliable and when it provides a provably better compute–quality tradeoff.
Q1:   Under what conditions is LoTTS guaranteed to outperform Global TTS?
A1:   As in Corollary 2, LoTTS is guaranteed to be better whenever localized updates are significantly cheaper than global ones ($C_{\ell}\ll C_{g}$). Every unit of compute allocated to LoTTS focuses exclusively on potentially defective regions, avoiding the large global harm term on (Ms)(M-s) clean patches. Thus even moderate precision and recall are sufficient for LoTTS to win.
Takeaway LoTTS becomes strictly more compute-efficient once its per-trial resampling cost is substantially lower than global resampling.
Q2:   In our implementation ($C_{\ell}\approx C_{g}$), why does LoTTS still win?
A2:   As in Corollary 3, when local and global trials cost the same, the comparison becomes purely statistical. In this case, the dominance condition in Corollary 1 becomes
$\rho\!\left[q\,\overline{\delta}-\Big(\tfrac{1}{\pi}-1\Big)h_{\ell}\,\overline{\gamma}\right]\;>\;\theta_{g}\,\overline{\delta}-\Big(\tfrac{M}{s}-1\Big)h_{g}\,\overline{\gamma}.$
LoTTS wins whenever its recall-weighted net benefit per selected patch exceeds the global benefit minus the massive global harm term.
Takeaway With equal compute, LoTTS still wins because its updates are localized, whereas the global harm term is too large to yield favorable net benefit.
Q3:   Why does the sparse-defect regime strongly favor LoTTS?
A3:   As shown by the sparse-regime analysis following Corollary 3, the sparse-defect regime strongly favors LoTTS because global resampling risks harming every clean patch. If $s\ll M$, the global-harm term
$\Big(\tfrac{M}{s}-1\Big)h_{g}\overline{\gamma}$
is extremely large. In contrast, LoTTS only touches $\rho s$ defective patches and $\rho s(1/\pi-1)$ mistakenly selected ones.
Takeaway Sparse defects strongly favor LoTTS, because local harm does not scale with the number of clean patches.
Q4:   How accurate must the mask be in avoiding clean patches? (precision)
A4:   As in equation 22, LoTTS only requires that each selected patch provides positive expected net gain:
$\pi>\Big(1+\tfrac{q\overline{\delta}}{h_{\ell}\overline{\gamma}}\Big)^{-1}.$
Takeaway Only modest precision is needed for selected patches to yield positive net gain.
Q5:   How many defective regions must the mask capture? (recall)
A5:   As in equation G.3.1, when $C_{\ell}\approx C_{g}$, LoTTS needs
$\rho\;>\;\frac{\theta_{g}\overline{\delta}-\big(\tfrac{M}{s}-1\big)h_{g}\overline{\gamma}}{q\overline{\delta}-(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}}.$
Takeaway LoTTS does not require perfect defect coverage to outperform global TTS.
Q6:   When might LoTTS fail to outperform global resampling?
A6:   As in Corollary 5, LoTTS may underperform only in regimes where its core advantages disappear:
  • defects are dense ($s\approx M$), eliminating the sparse-defect advantage;
  • precision is too low to avoid excessive harm;
  • local repair probability $q$ is dramatically worse than $\theta_{g}$.
Takeaway LoTTS underperforms only in corner cases where sparsity, mask precision, or local repair reliability breaks down.
Q7:   What is the practical bottom line?
A7:   In typical real-world scenarios, where defects are sparse and diffusion updates are locally benign, a mask with modest precision and recall already satisfies the dominance condition. Thus LoTTS provides a provably better compute–quality tradeoff than Global TTS, even when both operate under the same compute budget.
Takeaway In realistic settings, theoretical guarantees ensure LoTTS provides a superior compute–quality advantage.